### v1

# Semi-Bandit Learning for Monotone Stochastic Optimization

Abstract
Stochastic optimization is a widely used approach for optimization under uncertainty, where uncertain input parameters are modeled by random variables. Exact or approximation algorithms have been obtained for several fundamental problems in this area. However, a significant limitation of this approach is that it requires full knowledge of the underlying probability distributions. Can we still get good (approximation) algorithms if these distributions are unknown, and the algorithm needs to learn them through repeated interactions? In this paper, we resolve this question for a large class of “monotone” stochastic problems, by providing a generic online learning algorithm with $\sqrt{T\log T}$ regret relative to the best approximation algorithm (under known distributions). Importantly, our online algorithm works in a semi-bandit setting, where in each period, the algorithm only observes samples from the r.v.s that were actually probed. Our framework applies to several fundamental problems in stochastic optimization such as prophet inequality, Pandora’s box, stochastic knapsack, stochastic matchings and stochastic submodular optimization.
1 Introduction
Stochastic optimization problems have been a subject of intense investigation, as they offer a powerful lens through which we can handle uncertain inputs. In stochastic problems, uncertain input parameters are modeled by random variables, which are usually independent. Solutions (or policies) to a stochastic problem are sequential decision processes, where the next decision depends on all previously observed information. This approach has been applied to many domains: optimal stopping (Pandora’s box [Wei79, Sin18] and prophet inequality [SC84, KW12, FSZ21]), stochastic submodular optimization [AN16, GK17, INZ16], stochastic probing [GN13, ASW16, GNS17] and various other stochastic combinatorial optimization problems (e.g., knapsack [DHK16, DGV08] and matching [${}^{+}$, ${}^{+}$12]). A fundamental assumption in all these results is that the underlying probability distributions are known to the algorithm. While the known-distributions assumption holds in some settings, it is not satisfied in many practical applications, e.g., in the absence of historical data. Our main goal in this paper is to relax this assumption and handle stochastic problems with unknown distributions. We start by providing some concrete examples of the problems we consider.
Series testing.
There is a system with $n$ components, where each component $i$ is “working” independently with some probability $p_{i}$. All $n$ components must be working for the system to be functional. Moreover, it costs $c_{i}$ to test component $i$ and determine whether/not it is working. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Clearly, testing will continue until some component is found to be not working, or we have tested all components and found them to be working. In the standard setting, where the probabilities $\{p_{i}\}_{i=1}^{n}$ are known upfront, the greedy policy that tests components in decreasing order of $\frac{1-p_{i}}{c_{i}}$ is optimal [But72].
Pandora’s box.
There are $n$ items, where each item $i$ has an independent random value $\mathsf{X}_{i}$ with a known distribution $\mathcal{D}_{i}$. The realized value of $\mathsf{X}_{i}$ can be revealed by inspecting item $i$, which incurs cost $c_{i}$. The goal is to inspect a subset $S$ of the items (possibly adaptively) to maximize the expected utility $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. We emphasize that a solution to this problem can be quite complex: the choice of the next item to inspect may depend on the realizations of all previous items. Nevertheless, there is an elegant optimal solution to this problem when the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are known [Wei79].
Stochastic covering knapsack.
There are $n$ items, where each item $i$ has a (deterministic) cost $c_{i}$ and independent random reward $R_{i}$ with known distribution $\mathcal{D}_{i}$. The realized value of $R_{i}$ is only known when item $i$ is selected. Given a target $Q$, the goal is to select items sequentially (and adaptively) until the total reward is at least $Q$. The objective is to minimize the expected cost of selected items. Although the deterministic knapsack problem is already NP-hard, there is an elegant $3$-approximation algorithm for stochastic covering knapsack [DHK16] with known distributions.
In each of the above problems, what if the underlying probabilities/distributions are unknown? Is it still possible to obtain good performance guarantees? In order to address this question formally, we utilize the framework of online learning. Here, the algorithm interacts with an unknown-but-fixed input distribution $\mathcal{D}$ over multiple time periods. [^1] In the single period setting, it is easy to construct examples that rule out any reasonable performance bound for these problems with unknown distributions. In each period $t=1,\cdots T$, the online algorithm comes up with a solution/policy $\sigma^{t}$ to the stochastic problem and receives some feedback based on the performance of $\sigma^{t}$ on the (unknown) distribution $\mathcal{D}$.
The type of feedback received is a crucial component in this learning-based framework. The simplest setting is full feedback, where the algorithm receives one sample from every random variable (r.v.). However, this is unrealistic for stochastic problems because the policy $\sigma^{t}$ in period $t$ only observes the r.v.s corresponding to some subset of the items. For example, in series testing, $\sigma^{t}$ tests the $n$ components in some order until the first non-working component is found: if the first $k-1$ components are working and the $k^{th}$ component is not working then we would only observe the outcomes/samples of the first $k$ r.v.s (and the remaining $n-k$ r.v.s are unobserved). In this paper we consider the more natural setting of semi-bandit feedback, where in each period $t$, the algorithm only receives samples from the r.v.s that the policy $\sigma^{t}$ actually observed.
Our goal is to minimize the expected regret of the online algorithm, which is the difference between the total $T$-period objective of our algorithm and the optimum (which knows the distribution $\mathcal{D}$). Obtaining $o(T)$ regret dependence with respect to $T$ means that our online algorithm, which doesn’t know $\mathcal{D}$, asymptotically approaches the optimum. An additional difficulty arises from the fact that many problems that we consider are NP-hard (e.g., knapsack and submodular optimization). In such cases, we cannot hope to approach the optimum (via a polynomial algorithm). Here, we use the notion of $\alpha$-regret, where the online algorithm approaches the $\alpha$ approximately optimal value. In light of this, we ask: can we transform any offline (approximation) algorithm for a stochastic problem into an online learning algorithm that has low-regret with respect to this offline algorithm?
Our main result is an affirmative answer to this question. We give a general method for transforming any offline $\alpha$-approximation algorithm (with known distributions) to an online learning algorithm with $\alpha$-regret of $\widetilde{O}(\sqrt{T})$. It is well-known that the $\sqrt{T}$ regret bound cannot be improved, even in very special cases. Our method works for a wide range of stochastic problems that satisfy a natural monotonicity condition. This includes several fundamental problems in stochastic optimization such as prophet inequality, Pandora’s box, stochastic knapsack, stochastic matching, stochastic submodular maximization and submodular cover.
1.1 Problem Setup
We first set-up notation for stochastic optimization problems in the known distribution setting.
Stochastic Optimization.
Consider a stochastic problem $\mathcal{P}$ where the input consists of $n$ items with an independent random variable (r.v.) $\mathsf{X}_{i}$ associated with each item $i\in[n]=\{1,2,\cdots n\}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{X}_{i}$; that is, $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. We assume that each $\mathcal{D}_{i}$ has finite support, and denote the set of outcomes of all r.v.s by $O$. (We relax the discrete-distribution assumption in §2.6.) In order to determine the realization of any r.v. $\mathsf{X}_{i}$, we need to probe item $i$.
A solution or policy for $\mathcal{P}$ is given by a decision tree $\sigma$, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random realization of the probed item. Each node in decision tree $\sigma$ also corresponds to the current “state” of the policy, which is given by the sequence of previously-probed items along with their realizations; we will refer to nodes and states of the policy interchangeably. The root node of $\sigma$ is the starting state of the policy, at which point no item has been probed. Leaf nodes in $\sigma$ are also called terminal nodes/states. The policy execution under any realization $\mathbf{x}=(x_{1},\ldots,x_{n})$ corresponds to a root-leaf path $\sigma_{\mathbf{x}}$ in the decision tree $\sigma$, where at any node labeled by item $i$, path $\sigma_{\mathbf{x}}$ follows the branch corresponding to outcome $x_{i}$. Note that running policy $\sigma$ under $\mathbf{x}$ corresponds to traversing path $\sigma_{\mathbf{x}}$. We also define $S(\sigma,\mathbf{x})$ as the sequence of items probed by policy $\sigma$ under realization $\mathbf{x}$.
The cost of policy $\sigma$ depends on the realization $\mathbf{x}$, and we use $f(\sigma,\mathbf{x})\geq 0$ to denote the cost of policy $\sigma$ under realization $\mathbf{x}$. Specifically, $f(\sigma,\mathbf{x})$ depends only on the sequence of probed items $S(\sigma,\mathbf{x})$ and their realizations. We assume that cost is accrued only at terminal/leaf nodes: this is without loss of generality as every policy execution ends at a leaf node. [^2] In some problems, it is natural to accrue cost at each state that the policy goes through. We can easily convert such a cost structure to our setting by placing all the cost accrued from a policy execution at its terminal state. Note that we do not make any assumption on the cost function $f$: it does not have to be linear, submodular etc. Our goal is to find a policy $\sigma$ that minimizes the expected cost $f(\sigma):=\mathbb{E}_{\mathbf{x}}[f(\sigma,\mathbf{x})]$. We also allow for constraints in problem $\mathcal{P}$, which must be satisfied under all realizations (i.e., with probability $1$). Let $\mathcal{C}$ denote the set of feasible policies (that satisfy the constraints in $\mathcal{P}$). Sometimes, we work with randomized policies, where each node in the decision tree corresponds to a probability distribution over items (rather than a single item). Randomized policies are not any stronger than deterministic ones: there is always an optimal deterministic policy for this class of problems. Succinctly, the stochastic problem $\mathcal{P}$ is as follows (maximization problems are defined similarly).
$\text{minimize}\quad f(\sigma)=\mathbb{E}_{\mathbf{x}}\left[f(\sigma,\mathbf{x})\right]\quad\text{subject to}\quad\sigma\in\mathcal{C}.$
Observe that the number of nodes in policy $\sigma$ may be exponential. So, we are interested in “efficient” policies that can be implemented in polynomial-time for any realization $\mathbf{x}$. For some stochastic problems like series systems and Pandora’s box, efficient optimal policies are known. On the other hand, there are many problems like stochastic knapsack, matching and and set-cover, where optimal policies may not be efficient: in these cases, we will focus on (efficient) approximately optimal policies. We use $\sigma^{*}$ to denote an optimal policy, and denote its expected cost by $\mathtt{OPT}$. An $\alpha$-approximation algorithm for a stochastic problem $\mathcal{P}$ takes as input the probability distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ (along with any objective/constraint parameters) and returns a policy that has expected cost at most $\alpha\cdot\mathtt{OPT}$. We use the convention that for minimization problems, the approximation ratio $\alpha\geq 1$, whereas for maximization problems $\alpha\leq 1$.
The Online Semi-Bandit Setting.
In this setting, the distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ are unknown. Moreover, we have to repeatedly solve the stochastic problem $\mathcal{P}$ over $T$ time periods. We assume throughout the paper that $T\geq n$. The distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ remain the same across all $T$ periods. The goal is to simultaneously learn the distributions and converge to a good policy over time. At time $t\leq T$, we use all prior observations to present policy $\sigma^{t}$. The policies $\sigma^{t}$ may be different for each time $t\in[T]$. We measure our learning algorithm in terms of expected total regret. For minimization problems, we define it as follows (the definition for maximization problems is analogous).
$R(T)=\mathbb{E}_{\mathbf{x}^{1},\ldots,\mathbf{x}^{T}}\left[\sum_{t=1}^{T}\left(f(\sigma^{t},\mathbf{x}^{t})-f(\sigma^{*},\mathbf{x}^{t})\right)\right]=\mathbb{E}_{\mathbf{x}^{1},\ldots,\mathbf{x}^{T}}\left[\sum_{t=1}^{T}f(\sigma^{t},\mathbf{x}^{t})\right]-T\cdot\mathtt{OPT}.$ (1)
Here $\mathbf{x}^{t}$ represents the realization at time $t$. For conciseness, we use $\mathbf{h}^{t-1}$ to denote the history until time $t$; that is, $\mathbf{h}^{t-1}=(\mathbf{x}^{1},\ldots,\mathbf{x}^{t-1})$. Note that policy $\sigma^{t}$ is itself random because it depends on prior observations, i.e., on the history $\mathbf{h}^{t-1}$. The algorithm’s cost at time $t$ is $f(\sigma^{t},\mathbf{x}^{t})$, which depends additionally on the realizations $\mathbf{x}^{t}$ at time $t$. We can also re-write (1) as follows.
$R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})-f(\sigma^{*},\mathbf{x}^{t})\right]\right]=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})\right]-T\cdot\mathtt{OPT}.$ (2)
Note that many of the stochastic optimization problems are known to be NP-hard, even when the distributions are known. Thus, we do not expect to learn policies that approach $\mathtt{OPT}$. To get around this issue, we define the notion of $\alpha$-regret as follows.
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})\right]\right]-\alpha\cdot T\cdot\mathtt{OPT}.$ (3)
An important component in this learning setup is what type of feedback is received by the algorithm in each round. In the full feedback setting, the algorithm gets to know the entire sample $\mathbf{x}^{t}=(x_{1}^{t},\ldots,x_{n}^{t})$, which could be used to update beliefs regarding all the distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$. However, as mentioned earlier, this is unrealistic in our setting because in each period, the policy only observes a subset of r.v.s. So, we consider semi-bandit feedback, where the algorithm only observes the realizations of items that it probes. That is, at time $t$, the algorithm only observes $\left(X^{t}_{i}:i\in S(\sigma^{t},\mathbf{x}^{t})\right)$ where $S(\sigma^{t},\mathbf{x}^{t})$ is the set of items probed by policy $\sigma^{t}$ before it terminates.
1.2 Results and Techniques
For any stochastic problem instance $\mathcal{I}$ and distribution $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$, let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ denote the optimal cost of instance $\mathcal{I}$ with r.v.s having distribution $\mathbf{{D}}$. We first define a monotonicity condition for a stochastic problem $\mathcal{P}$ which will be used in our main result.
Definition 1.1 (Monotonicity).
A stochastic problem $\mathcal{P}$with a minimization objective is up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$. Similarly, the problem is said to be down-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$. The definitions for maximization problems are analogous.
Theorem 1.1.
Suppose that a stochastic problem $\mathcal{P}$ has an $\alpha$-approximation algorithm, and it is either up-monotone or down-monotone. Then, there is a polynomial time semi-bandit learning algorithm for $\mathcal{P}$ (with unknown distributions) that has $\alpha$-regret $O(nkf_{max}\sqrt{T\log(kT)})$. Here, $n$ is the number of items, $k$ is the maximum support size of any distribution, $f_{max}$ is the maximal value of the objective function $f$ (over all realizations) and $T$ is the number of time periods.
We note that the dependence on the support-size $k$ can be replaced by the maximum number of “thresholds” used at any node of the stochastic policy (which corresponds to the number of branches out of the node). For most of our applications, we have approximate policies with $O(1)$ thresholds: so the regret bound is just $O(nf_{max}\sqrt{T\log T})$. See Theorem 2.9 for the formal statement, which also handles continuous r.v.s.
Consequently, we get polynomial time online learning algorithms for a number of stochastic optimization problems. We list a few of them next, highlighting the dependence on $n$ and $T$ (see §3 for further details and more applications).
  • Pandora’s box: $1$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Prophet inequality: $1$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Matroid Prophet inequality [KW19]: $\frac{1}{2}$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Stochastic (maximization) knapsack [DGV08, Ma18]: $\frac{1}{2}$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Stochastic covering knapsack [DHK16]: $3$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
The first two results improve over the regret bounds of $O\left(n^{3}\sqrt{T}\log(T)\right)$ and $O\left(n^{4.5}\sqrt{T}\log(T)\right)$ for prophet inequality and Pandora’s box from [GKSW24]. To the best of our knowledge, our results for all other applications are the first $\alpha$-regret bounds of $\sqrt{T}$. We note that the result of [GKSW24] for Pandora’s box and Prophet inequality also holds under the more restrictive bandit feedback model (only the objective value is observed). However, it is limited to these two problems, whereas our approach for semi-bandit feedback provides $\sqrt{T}$ regret for a much broader class of problems. Prior work (based on sample complexity) only gives $T^{2/3}$ regret for these problems [GHTZ21].
The regret bound in Theorem 1.1 is nearly optimal in the following sense. There is an $\Omega(\sqrt{T})$ lower bound on the regret even in the much simpler setting of multi-armed-bandits [LR85]. Furthermore, if we move beyond our setting of independent-identically-distributed (i.i.d.) distributions across periods, to the “adversarial” setting with different distributions for each period, then there is a linear $\Omega(T)$ lower-bound on regret (even for prophet inequality) [GKSW24].
Technical Overview.
At a high-level, our algorithm is based on the principle of optimism in the face of uncertainty which is well-studied in the multi-armed bandits literature: see e.g., the UCB algorithm [LS20]. Given observations from previous rounds, our algorithm first constructs a modified empirical distribution $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ which stochastically dominates the true distribution $\mathbf{{D}}$. It then executes the policy $\sigma$ given by the offline (approximation) algorithm under distribution $\mathbf{{E}}$.
We now discuss how to bound the regret of the policy $\sigma$ under the true distribution $\mathbf{{D}}$. To keep things simple here, we assume that $\alpha=1$ and the problem is up-monotone. The difficulty in our setting is that the decision to probe/observe an item is random: it depends on the choice of policy $\sigma$ and the underlying distribution $\mathbf{D}$ of items. In contrast, the classical multi-armed bandits setting allows direct control to the algorithm on which item to probe. This gives algorithms such as UCB a direct control on the rate of exploration of individual items and allows them to bound the regret of individual items in terms of this rate. How do we bound the rate of exploration of different items in our setting when we do not have direct control on which items are probed? Moreover, how do we bound the regret contribution of individual items in terms of their rate of exploration? The key insight in our analysis is the following “stability” bound, which answers the above questions. Given product distributions $\mathbf{{D}}$ and $\mathbf{{E}}$, where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we show [^3] For policy $\sigma$ and distribution $\mathbf{D}$, $f(\sigma|\mathbf{{D}})$ is the expected objective of $\sigma$ when the r.v.s have distribution $\mathbf{D}$.
$f(\sigma|\mathbf{{D}})-f(\sigma^{*}|\mathbf{{D}})\,\,\leq\,\,f(\sigma|\mathbf{{D}})-f(\sigma|\mathbf{{E}})\,\,\leq\,\,f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma)\cdot\epsilon_{i}\,,$
where $\sigma^{*}$ (resp. $\sigma$) is an optimal policy under $\mathbf{{D}}$ (resp. $\mathbf{{E}}$), $Q_{i}(\sigma)$ is the probability that policy $\sigma$ probes item $i$, and $\epsilon_{i}$ is the total-variation distance between $\mathcal{E}_{i}$ and $\mathcal{D}_{i}$. The first inequality above follows from the monotone property. The second inequality requires analyzing the decision tree of $\sigma$ carefully and is a key technical contribution of this paper (see §2.2).
The above stability bound can be interpreted as follows: the contribution of item $i$ in the total regret is the probability that it is probed times the total variation error in estimating $\mathcal{D}_{i}$. The dependence on the probability of probing is crucial here. This allows us to analyze the regret using a charging argument where we charge item $i$ for regret in a pay-per-use manner. The paths in the execution tree of the online algorithm that probe $i$ will pay for the error in estimation of $\mathcal{D}_{i}$, but then $\epsilon_{i}$ will also reduce along these paths. Moreover, we can bound the overall regret irrespective of the policy $\sigma$ that is used at periods. The error $\epsilon_{i}\approx\sqrt{1/m}$ when $i$ has been probed $m$ times in the past, which gives a total regret contribution of $\sum_{m=1}^{T}\sqrt{1/m}=O(\sqrt{T})$ for any item. The details appear in §2.1. Finally, we lose an extra logarithmic factor to ensure that our empirical distribution $\mathbf{E}$ stochastically dominates $\mathbf{D}$ with high probability.
Let us contrast our algorithm (and stability bound) with another algorithm that balances the explore-exploit trade-off in a different manner: the explore-then-commit algorithm. This algorithm will explore each item $O(T^{2/3})$ times (say, by probing it first in policy $\sigma$) and then commit to a policy based on these $O(nT^{2/3})$ observations. Using the sample complexity bound of [GHTZ21] it is easy to show that this algorithm achieves $O(T^{2/3})$ regret; see §A for details. However, each item is explored an equal number of times regardless of its cost and we might end up exploring some high-cost spurious items too many times. Our stability bound highlights a crucial difference between our algorithm and this algorithm: we only explore items that are needed in the policy. If there are items which are rarely used in the policy, they will only be explored infrequently and will not contribute to the regret as well. This also explains why our approach achieves a near-optimal $\sqrt{T\log T}$ regret bound compared to the $T^{2/3}$ bound of the sampling-based approach.
1.3 Related Work
As mentioned earlier, stochastic optimization problems (under known distributions) have been studied extensively. Several papers have extended the classic prophet inequality [SC84] and Pandora’s box [Wei79] results to more complex settings, e.g., [KW12, RS17, KWW16, Sin18, FSZ21, BFLL23]. Moreover, good approximation bounds have been achieved for stochastic versions of various combinatorial optimization problems [DGV08, GK17, Ma18, GKNR15, INZ16, GV06, GN13, GNS17, DHK16, ${}^{+}$, ${}^{+}$12].
There has also been extensive work in online learning, where one considers unknown distributions (in the stochastic learning setting), see e.g. books [CBL06, LS20, Haz22]. All our stochastic problems can be modeled as Markov decision processes (MDPs) with unknown transition probabilities, and there have been some works on achieving sublinear regret in this setting [AOM17, JABJ18, ${}^{+}$20]. However, the regret bounds from these work have a polynomial dependence on the “state space” of the MDP, which in our setting, is exponential in $n$ (the number of r.v.s). It is also known that any online algorithm for arbitrarty MDPs incurs an $\Omega(\sqrt{S\cdot T})$ regret, where $S$ is the size of the state space [JOA10, OR16]. In contrast, we obtain $poly(n)\cdot\sqrt{T\log T}$ regret.
Although both stochastic optimization and online learning have been subjects of comprehensive research, there has been relatively limited work at the intersection of these two domains. For instance, [${}^{+}$13] studied a special case of adaptive submodular maximization [GK17] in the bandit setting. Some recent papers [${}^{+}$20, GT22] considered online learning for Pandora’s box in a correlated setting, which is much harder than the independent case: one needs to resort to weaker “partially adaptive” benchmarks here. [EHLM19] also considered a different “multi armed” Pandora’s box problem, and obtained a $\frac{1}{2}$-competitive ratio relative to a suitable benchmark. Drawing closer to our work is [GKSW24], which explores Pandora’s box and prophet inequality within the online learning framework with bandit feedback, which is even more restrictive than our semi-bandit feedback. In bandit feedback, the algorithm only observes the realized objective value of its policy $\sigma$. [GKSW24] gave algorithms with $1$-regret $O\left(n^{3}\sqrt{T}\log(T)\right)$ and $O\left(n^{4.5}\sqrt{T}\log(T)\right)$ for Pandora’s box and prophet inequality respectively. While our feedback model is not as general as [GKSW24], we think that semi-bandit feedback already captures the issue of partial observations in policies. Moreover, under semi-bandit feedback we obtain a general framework for online learning algorithms, that applies to several stochastic optimization problems.
There have also been several works studying sample complexity bounds for several stochastic optimization problems. The goal in these works is to understand how much data is necessary and sufficient to guarantee near-optimal algorithms. Such results have been obtained for single-parameter revenue maximization [CR14, RS16] and prophet inequality [CDFS19]. Recent work [GHTZ21] gives optimal sample complexity bounds for stochastic optimization problems that exhibit strong monotonicity, which is a slightly stronger condition than our monotonicity definition. These results imply $\sqrt{T}$ regret under the full-feedback model, but only ${T}^{2/3}$ regret under the semi-bandit feedback model considered in this paper (see Appendix A).
Our results are related in spirit to some previous online-to-offline reductions. In particular, [KV05] considered combinatorial optimization with linear objective functions, and obtained $1$-regret of $O(n\sqrt{T})$ under full-feedback (assuming an exact offline algorithm). Then, [KKL09] extended this result to linear problems with only an $\alpha$-approximation algorithm, and obtained $\alpha$-regret of $O(n\sqrt{T})$ under full-feedback and $O(nT^{2/3})$ under bandit-feedback. Recent work of [${}^{+}$21] considers certain combinatorial optimization problems with non-linear objectives, assuming an $\alpha$-approximation via a greedy-type algorithm. For such problems, [${}^{+}$21] obtained $\sqrt{T}$ regret under full-feedback and $T^{2/3}$ regret under bandit-feedback. We note that all these results work in the (harder) adversarial online setting, rather than our i.i.d. online setting. However, our result handles a much wider class of problems, with arbitrary objectives. Moreover, as noted before, there is an $\Omega(T)$ regret lower-bound for many of our applications in the adversarial setting.
1.4 Preliminaries
We present some preliminaries that are required in our algorithm and proofs.
Definition 1.2 (Stochastic Dominance).
A probability distribution $\mathcal{E}$ (over $\mathbb{R}$) stochastically dominates another distribution $\mathcal{D}$ if, for all $a\in\mathbb{R}$, we have: $\mathbf{P}_{X\leftarrow\mathcal{E}}(X\geq a)\geq\mathbf{P}_{Y\leftarrow\mathcal{D}}(Y\geq a)$.
We use $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ to denote that distribution $\mathcal{E}$ stochastically dominates $\mathcal{D}$. For product distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$, we say that $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$ if $\mathcal{E}_{i}$ stochastically dominates $\mathcal{D}_{i}$ for all $i\in[n]$; we also denote this by $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$.
Probability Distribution Metric.
Let $\mathtt{TV}(\mathcal{D},\mathcal{E})$ denote the total variation distance between discrete distributions $\mathcal{D}$ and $\mathcal{E}$. The total variation distance is half of the $\ell_{1}$ distance between the two distributions, i.e., $\mathtt{TV}(\mathcal{D},\mathcal{E})=\frac{1}{2}\cdot||\mathcal{D}-\mathcal{E}||_{1}$. The following standard result (see, for example, Lemma B.8 in[GVdV17]) bounds the total variation distance between product distributions.
Lemma 1.2.
Given product distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ over $n$ r.v.s, we have
$\mathtt{TV}(\mathbf{{D}},\mathbf{{E}})\leq\sum_{i\in[n]}\mathtt{TV}(\mathcal{D}_{i},\mathcal{E}_{i}).$
Consider independent r.v.s $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$ where $\mathsf{T}_{i}$ denotes the domain of $\mathsf{X}_{i}$. Let $h$ be a function from $\mathbf{{T}}=\mathsf{T}_{1}\times\cdots\times\mathsf{T}_{n}$ to $[0,U]$; that is, $h$ is a function on the outcomes of the random variables that is bounded by $U$. Thus, for any $\mathbf{x}\in\mathbf{{T}}$, $h(\mathbf{x})$ denotes the value of $h$ on the outcome $\mathbf{x}=(x_{1},\ldots x_{n})$. Given a product distribution $\mathbf{{P}}$ over the r.v.s, define
$h(\mathbf{{P}})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{P}}}\left[h(\mathbf{x})\right].$
The following useful fact bounds the difference in function value at two different distributions.
Lemma 1.3.
Given discrete distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ over $n$ random variables $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$, we have
$|h(\mathbf{{D}})-h(\mathbf{{E}})|\,\,\leq\,\,U\cdot\mathtt{TV}(\mathbf{{D}},\mathbf{{E}}).$
2 The Online Framework
In this section we present our main algorithm. We will assume access to an $\alpha$-approximation algorithm $\mathtt{ALG}$ for the stochastic problem $\mathcal{P}$. For concreteness, we assume that $\mathcal{P}$ is a minimization problem. Thus, given an instance of $\mathcal{P}$ (with a probability distribution for each item), $\mathtt{ALG}$ finds a policy of expected cost at most $\alpha$ times the optimum. We also assume that $\mathcal{P}$ is up-monotone; see Definition 1.1. The online framework for down-monotone problems and maximization problems is almost identical: the changes are explained in §2.4 and §2.5.
Our online algorithm is based on the principle of optimism in the face of uncertainty, and is very simple to describe. At each time step $t$, we construct an “optimistic” empirical distribution $\mathbf{{E}}^{t}$ that stochastically dominates the true (unknown) distribution $\mathbf{{D}}$. Then, we run the offline algorithm $\mathtt{ALG}$ on this empirical distribution $\mathbf{{E}}^{t}$ to obtain policy $\sigma^{t}$, which is the online policy at time $t$.
Computing stochastically dominating distributions.
Given i.i.d. samples of any random variable, we show that it is possible to compute a stochastically dominating empirical distribution that has a small total-variation distance from the true distribution.
Theorem 2.1.
There is an algorithm $\mathtt{EmpStocDom}$ that, given $m$ i.i.d. samples from a distribution $\mathcal{D}$ with finite support-size $k$, and parameter $\delta>0$, computes a distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$:
  • $\mathcal{E}$ stochastically dominates $\mathcal{D}$, and
  • the total-variation distance $\mathtt{TV}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2k/\delta)}{2m}}$.
The algorithm $\mathtt{EmpStocDom}$ and proof of this result are deferred to §2.3. The main idea is to “shift” some probability mass in the usual empirical distribution from low to high values. Algorithm 1 describes our online framework, which uses algorithm $\mathtt{EmpStocDom}$.
Algorithm 1 Online-to-Offline framework
1:for $t=1,\ldots T$ do
2:     for each item $i\in[n]$, obtain distribution ${\mathcal{E}}^{t}_{i}$ by running $\mathtt{EmpStocDom}$ with $\delta=\frac{2}{(nT)^{3}}$ on all samples of r.v. $\mathsf{X}_{i}$ observed so far.
3:     obtain policy $\sigma^{t}$ by running the offline algorithm $\mathtt{ALG}$ on distribution $\mathbf{{E}}^{t}=\left\{{\mathcal{E}}^{t}_{i}\right\}_{i=1}^{n}$.
4:     run policy $\sigma^{t}$ on realization $\mathbf{x}^{t}$ and observe the probed items $S(\sigma^{t},\mathbf{x}^{t})$.
We now analyze Algorithm 1, and prove Theorem 1.1. The basic idea in the analysis is that as we get more and more samples with increasing $t$, the total variation distance between $\mathbf{{E}}^{t}$ and $\mathbf{{D}}$ reduces, and the policy $\mathtt{ALG}(\mathbf{{E}}^{t})$ will become closer and closer to the optimal policy.
We will assume, without loss of generality, that we have observed at least one sample from each random variable. This can be ensured by adding $n$ special time periods and choosing for each time $i\in[n]$, a policy $\sigma^{i}$ that first probes item $i$; note that this contributes at most $n\cdot f_{max}$ to the total regret. Let $N_{j}^{t}$ be the number of times r.v. $\mathsf{X}_{j}$ has been sampled before time $t$. By our assumption, $N_{j}^{1}=1$ for all $j\in[n]$. Note that $N_{j}^{t}$ is a random variable and depends on the history $\mathbf{h}^{t-1}=\left(\mathbf{x}^{1},\cdots\mathbf{x}^{t-1}\right)$. When needed, we will use $N_{j}^{t}(\mathbf{h}^{t-1})$ to make this dependence explicit. The following result uses Theorem 2.1 and union bound.
Lemma 2.2.
With probability at least $1-\frac{1}{nT}$, we have
$\mathcal{E}_{j}^{t}\text{ stochastically dominates }\mathcal{D}_{j}\text{ and }\mathtt{TV}\left(\mathcal{E}^{t}_{j},\mathcal{D}_{j}\right)<k\cdot\sqrt{\frac{3\log(knT)}{2N_{j}^{t}}},\,\,\mbox{ for all }j\in[n]\mbox{ and }t\in[T].$
Proof.
For any $j\in[n]$ and $t\in[T]$, let $B_{j}^{t}$ denote the event that the stated condition does not hold for $j$ and $t$. Using Theorem 2.1 with $\delta=2/(nT)^{3}$, we obtain for any $j\in[n]$ and $t\in[T]$,
$\mathbf{P}\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\delta,\qquad\forall m\geq 1.$
So, for any $j$ and $t$, we have by union bound,
$\mathbf{P}(B_{j}^{t})\leq\sum_{m=1}^{T}\mathbf{P}\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\sum_{m=1}^{T}\delta=\delta T.$
Again, by union bound,
$\mathbf{P}\left(\vee_{j=1}^{n}\vee_{t=1}^{T}B_{j}^{t}\right)\leq\sum_{j=1}^{n}\sum_{t=1}^{T}\mathbf{P}(B_{j}^{t})\leq nT^{2}\delta\leq\frac{1}{nT}.$
The last inequality uses $n\geq 2$. This completes the proof. ∎
We now define a good event $G$ that corresponds to the condition in Lemma 2.2 holding true (for all $j\in[n]$ and $t\in[T]$). Next, we complete the proof assuming that $G$ holds. At the end, we show how to remove this assumption (using the fact that $G$ holds with high probability).
We now state a crucial “stability” property for the stochastic problem $\mathcal{P}$.
Lemma 2.3 (Stability lemma).
Consider a stochastic problem that is up-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
$f(\sigma)-\alpha\cdot f(\sigma^{*})=\mathbb{E}_{\mathbf{x}\sim\mathbf{D}}\left[f(\sigma,\mathbf{x})-\alpha\cdot f(\sigma^{*},\mathbf{x})\right]\leq f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma)\cdot\epsilon_{i}\,,$ (4)
where ${Q}_{i}(\sigma)$ denotes the probability that item $i$ is probed by policy $\sigma$ under distribution $\mathbf{{D}}$.
This lemma relies on the monotonicity assumption, and is proved in §2.2.
2.1 Completing the Proof of Theorem 1.1
We now complete the proof of Theorem 1.1 using Lemma 2.3.
Bounding the regret as sum over time $t$.
To bound the overall regret, it suffices to bound the regret at each time $t\in[T]$, defined as:
$R^{t}(\mathbf{h}^{t-1}):=\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})-\alpha\cdot f(\sigma^{*},\mathbf{x}^{t})\right]=f(\sigma^{t})-\alpha\cdot f(\sigma^{*}),$
where $\mathbf{h}^{t-1}$ is the history at time $t$, policy $\sigma^{t}=\mathtt{ALG}(\mathbf{E}^{t})$ and policy $\sigma^{*}$ is the optimal policy for the stochastic problem instance (under the true distribution $\mathbf{D}$), and policy $\sigma^{t}$ is the policy used by our online algorithm at time $t$. By (3), the overall regret is
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[R^{t}(\mathbf{h}^{t-1})\right].$
For each time $t\in[T]$ and history $\mathbf{h}^{t-1}$, we apply Lemma 2.3 with distributions $\mathbf{E}^{t}(\mathbf{h}^{t-1})$ and $\mathbf{D}$, and parameters $\epsilon_{i}^{t}(\mathbf{h}^{t-1})=k\cdot\sqrt{\frac{3\log(knT)}{2N_{i}^{t}(\mathbf{h}^{t-1})}}$ for all $i\in[n]$. Note that $\mathtt{TV}(\mathcal{E}^{t}_{i},\mathcal{D}_{i})\leq\epsilon_{i}^{t}(\mathbf{h}^{t-1})$ and $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$ for each $i\in[n]$ because we assumed the good event $G$. Hence, Lemma 2.3 implies that:
$R^{t}(\mathbf{h}^{t-1})\leq f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma^{t})\cdot\epsilon_{i}^{t}(\mathbf{h}^{t-1}).$ (5)
We note that $R^{t}$, $\sigma^{t}$, $N^{t}$ and $\epsilon^{t}$ all depend on the history $\mathbf{h}^{t-1}$; to keep notation simple, we drop the explicit dependence on $\mathbf{h}^{t-1}$. Combining the above two inequalities,
$\alpha\text{-}R(T)\leq f_{\max}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}{Q}_{i}(\sigma^{t})\cdot\epsilon_{i}^{t}\right]=kf_{\max}\sqrt{1.5\log(knT)}\cdot\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\frac{{Q}_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right].$
It now suffices to show
$\sum_{t=1}^{T}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{{Q}_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right]\leq 2n\sqrt{T}.$ (6)
Indeed, combining (6) with the above bound on regret, we get $\alpha\text{-}R(T)\leq knf_{max}\sqrt{6\log(knT)}$, which completes the proof of Theorem 1.1.
Proving (6) as sum over decision paths.
We refer to the full history $\mathbf{h}^{T}=\left(\mathbf{x}^{1},\cdots,\mathbf{x}^{T}\right)$ of the algorithm as its decision path. The main idea in this proof is to view the left-hand-side in (6) as a sum over decision paths rather than a sum over time. To this end, define:
$Z_{i}(\mathbf{h}^{T}):=\sum_{t=1}^{T}\frac{\mathbf{I}\left[\text{item $i$ probed by policy }\sigma^{t}(\mathbf{h}^{t-1})\right]}{\sqrt{N^{t}_{i}(\mathbf{h}^{t-1})}},\quad\forall i\in[n].$ (7)
Above, $\mathbf{I}$ is the indicator function. Also, let $Z(\mathbf{h}^{T}):=\sum_{i=1}^{n}Z_{i}(\mathbf{h}^{T})$. By linearity of expectation,
$\mathbb{E}_{\mathbf{h}^{T}}\left[Z_{i}(\mathbf{h}^{T})\right]$ $=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{T}}\left[\frac{\mathbf{I}\left[i\text{ probed by }\sigma^{t}(\mathbf{h}^{t-1})\right]}{\sqrt{N^{t}_{i}(\mathbf{h}^{t-1})}}\right]\,\,=\,\,\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1},\mathbf{x}^{t}}\left[\frac{\mathbf{I}\left[i\text{ probed by }\sigma^{t}(\mathbf{h}^{t-1})\right]}{\sqrt{N^{t}_{i}(\mathbf{h}^{t-1})}}\right]$
$=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{1}{\sqrt{N^{t}_{i}}}\cdot\mathbf{P}_{\mathbf{x}^{t}}[i\text{ probed by }\sigma^{t}]\right]\,\,=\,\,\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{Q_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right],$ (8)
where the second equality uses the fact that event {$i$ probed by $\sigma^{t}$} only depends on $\mathbf{h}^{t}=(\mathbf{h}^{t-1},\mathbf{x}^{t})$, the third equality uses the fact that $N^{t}_{i}$ only depends on $\mathbf{h}^{t-1}$ and that $\mathbf{x}^{t}$ is independent of $\mathbf{h}^{t-1}$, and the last equality is by the definition of $Q_{i}(\sigma^{t})$ and the fact that $\mathbf{x}^{t}\sim\mathbf{D}$. Using (8) and adding over $i\in[n]$, we get
$\sum_{t=1}^{T}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{{Q}_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right]=\sum_{i=1}^{n}\mathbb{E}_{\mathbf{h}^{T}}\left[Z_{i}(\mathbf{h}^{T})\right]=\mathbb{E}_{\mathbf{h}^{T}}[Z(\mathbf{h}^{T})].$
Therefore, proving (6) is equivalent to:
$\mathbb{E}_{\mathbf{h}^{T}}[Z(\mathbf{h}^{T})]=\sum_{i=1}^{n}\mathbb{E}_{\mathbf{h}^{T}}[Z_{i}(\mathbf{h}^{T})]\leq 2n\sqrt{T}$ (9)
We now prove (9) by showing that $\mathbb{E}_{\mathbf{h}^{T}}[Z_{i}(\mathbf{h}^{T})]\leq 2\sqrt{T}$ for each $i\in[n]$. Indeed,
$Z_{i}(\mathbf{h}^{T})=\sum_{t=1}^{T}\frac{\mathbf{I}\left[i\text{ probed by }\sigma^{t}(\mathbf{h}^{t-1})\right]}{\sqrt{N^{t}_{i}(\mathbf{h}^{t-1})}}\leq\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}.$
The first inequality uses the fact that $N^{t}_{i}(\mathbf{h}^{t-1})$ equals the number of probes of item $i$ in the first $t-1$ time steps: so $N^{t+1}_{i}=N^{t}_{i}+1$ whenever item $i$ is probed by $\sigma^{t}(\mathbf{h}^{t-1})$. This completes the proof of (9) and hence (6).
Removing the “good” event assumption.
In the analysis above, we assumed that event $G$ holds in (5). We now modify (5) as follows (which holds irrespective of $G$).
$R^{t}(\mathbf{h}^{t-1})\leq kf_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma^{t})\cdot\epsilon_{i}^{t}(\mathbf{h}^{t-1})+f_{max}\cdot\mathbf{I}[\overline{G}].$
We used the fact that the maximum 1-step regret is $f_{max}$. Combined with the previous analysis (which handles the first term), we have
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}[R^{t}(\mathbf{h}^{t-1})]\leq knf_{max}\sqrt{6\log(knT)}\,+\,Tf_{max}\cdot\mathbf{P}[\overline{G}]\leq knf_{max}\sqrt{6\log(knT)}+f_{max}.$
This completes the proof of Theorem 1.1 for up-monotone problems.
2.2 Proving the Stability Lemma
We now prove Lemma 2.3. Recall that there are two product distributions $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. Let $\sigma=\mathtt{ALG}(\mathbf{E})$ be the policy returned by the (offline) $\alpha$-approximation algorithm, and $\sigma^{*}$ be an optimal policy under distribution $\mathbf{D}$. We want to upper bound $f(\sigma)-\alpha\cdot f(\sigma^{*})$. Let $\mathcal{C}$ denote the set of all feasible policies to the given instance. Then,
$\mathtt{OPT}(\mathbf{D})=\min_{\tau\in\mathcal{C}}\mathbb{E}_{\mathbf{x}\sim\mathbf{D}}[f(\tau,\mathbf{x})]\quad\text{and}\quad\mathtt{OPT}(\mathbf{E})=\min_{\tau\in\mathcal{C}}\mathbb{E}_{\mathbf{x}\sim\mathbf{E}}[f(\tau,\mathbf{x})].$
We use $f(\tau|\mathbf{{U}}):=\mathbb{E}_{\mathbf{x}\sim\mathbf{{U}}}[f(\tau,\mathbf{x})]$ to denote the expected cost of any policy $\tau$ when the r.v.s have product distribution $\mathbf{{U}}$. Note that $f(\tau)=f(\tau|\mathbf{D})$ where $\mathbf{D}$ is the true distribution. We now have:
$f(\sigma)-\alpha\cdot f(\sigma^{*})$ $=f(\sigma|\mathbf{D})-\alpha\cdot f(\sigma^{*}|\mathbf{D})=f(\sigma|\mathbf{D})-f(\sigma|\mathbf{E})+f(\sigma|\mathbf{E})-\alpha\cdot f(\sigma^{*}|\mathbf{D})$
$\leq f(\sigma|\mathbf{D})-f(\sigma|\mathbf{E})+\alpha\cdot\mathtt{OPT}(\mathbf{E})-\alpha\cdot f(\sigma^{*}|\mathbf{D})$ (10)
$=f(\sigma|\mathbf{D})-f(\sigma|\mathbf{E})+\alpha\cdot\mathtt{OPT}(\mathbf{E})-\alpha\cdot\mathtt{OPT}(\mathbf{D})\,\,\leq\,\,f(\sigma|\mathbf{D})-f(\sigma|\mathbf{E}).$ (11)
Inequality (10) uses the fact that $\sigma$ is an $\alpha$-approximate policy to the instance with distribution $\mathbf{E}$. In (11), the equality uses the fact that $\sigma^{*}$ is an optimal policy for the instance with distribution $\mathbf{D}$, and the inequality is by up-monotonicity (Definition 1.1) and $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$.
Equivalent view of adaptive policy $\sigma$.
Let $N$ denote the number of nodes in decision tree $\sigma$ (this may be exponential, but it is only used in the analysis). Note that there is a partial ordering of the nodes of $\sigma$ based on ancestor-descendent relationships in the tree. We index the nodes in $\sigma$ from $1$ to $N$ according to this partial order, so that $u<v$ for any node $v$ that is a child of node $u$. Recall that each node $v\in\sigma$ is labeled by one of the $n$ items. Note that the same item $i\in[n]$ may label multiple nodes of $\sigma$; however, in any policy execution (i.e., root-leaf path in $\sigma$) we will encounter at most one node labeled by item $i$. Hence, we can equivalently view policy $\sigma$ as having an item $\mathsf{X}_{v}$ with independent distribution $\mathcal{D}_{v}$ at each node $v\in\sigma$. (This involves making several independent copies of each item, which does not affect the policy execution as at most one copy of each item is seen on any root-leaf path.)
Bounding $|f({\sigma}|\mathbf{{D}})-f({\sigma}|\mathbf{{E}})|$.
Based on the above view of $\sigma$, we use $\mathcal{D}_{v}$ (resp. $\mathcal{E}_{v}$) to denote the independent distribution at each node $v\in\sigma$ under the joint distribution $\mathbf{D}$ (resp. $\mathbf{E}$). Using the above indexing of the nodes in $\sigma$, we define the following hybrid product distributions: for each $v\in[N]$ let $\mathbf{{H}}^{v}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{v}\times\mathcal{E}_{v+1}\times\cdots\times\mathcal{E}_{N}$. Observe that $\mathbf{{H}}^{N}=\mathbf{{D}}$ and $\mathbf{{H}}^{0}=\mathbf{{E}}$. Using a telescoping sum, we can write:
$f(\sigma\mid\mathbf{D})-f(\sigma\mid\mathbf{E})=\sum_{v=1}^{N}f(\sigma\mid\mathbf{{H}}^{v})-f(\sigma\mid\mathbf{{H}}^{v-1}).$
Crucially, we now show that for every node $v\in[N]$,
$\mid f(\sigma\mid\mathbf{{H}}^{v})-f(\sigma\mid\mathbf{{H}}^{v-1})\mid\,\,\leq\,\,f_{\max}\cdot Q_{v}(\sigma)\cdot\epsilon_{v},$ (12)
where $Q_{v}(\sigma)$ is the probability that $\sigma$ reaches node $v$ under distribution $\mathbf{D}$ and $\epsilon_{v}=\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v})$. We first complete the proof of the lemma using (12). Note that for any node $v$ labeled by item $i\in[n]$, we have $\epsilon_{v}=\mathtt{TV}(\mathcal{D}_{i},\mathcal{E}_{i})\leq\epsilon_{i}$.
$\big{|}f(\sigma\mid\mathbf{D})-f(\sigma\mid\mathbf{E})\big{|}$ $\leq\sum_{v=1}^{N}\big{|}f(\sigma\mid\mathbf{{H}}^{v})-f(\sigma\mid\mathbf{{H}}^{v-1})\big{|}\leq f_{max}\cdot\sum_{v=1}^{N}Q_{v}(\sigma)\cdot\epsilon_{v}$
$=f_{max}\cdot\sum_{i=1}^{n}\epsilon_{i}\sum_{v\in[N]:\mbox{ labeled }i}Q_{v}(\sigma)\,\,=\,\,f_{max}\cdot\sum_{i=1}^{n}\epsilon_{i}\cdot Q_{i}(\sigma).$ (13)
The last equality uses the fact that the probability $Q_{i}(\sigma)$ of probing $i$ equals the probability of reaching some node $v\in[N]$ labeled $i$.
Towards proving (12), we introduce the following notation. Let $Q_{v}(\sigma\mid\mathbf{{U}})$ denote the probability that $\sigma$ reaches node $v$ under any distribution $\mathbf{{U}}$. Note that the event “$\sigma$ reaches node $v$” only depends on the random realizations at the ancestor nodes of $v$, which are all contained in $\{u\in[N]:u<v\}$ (by our indexing of nodes). Hence,
$Q_{v}(\sigma\mid\mathbf{{H}}^{v})=Q_{v}(\sigma\mid\mathbf{{H}}^{v-1})=Q_{v}(\sigma\mid\mathbf{{D}})=Q_{v}(\sigma),$ (14)
because each node $\{u\in[N]:u<v\}$ has the same distribution ($\mathcal{D}_{u}$) under $\mathbf{D}$, $\mathbf{{H}}^{v}$ and $\mathbf{{H}}^{v-1}$.
Below, let ${\cal R}_{v}$ denote the event that $\sigma$ reaches node $v$. Using the fact that each node $w\in[N]\setminus\{v\}$ has the same distribution (either $\mathcal{D}_{w}$ or $\mathcal{E}_{w}$) under both $\mathbf{{H}}^{v}$ and $\mathbf{{H}}^{v-1}$, we get:
[…]
We now bound the difference in expectation conditional on event ${\cal R}_{v}$. Let $\rho$ denote the subtree of $\sigma$ rooted at $v$ (including $v$ itself). We also use $\rho$ to denote the set of nodes in this subtree. Let $L$ denote all leaf nodes in subtree $\rho$, and for each $\ell\in L$ let $f_{\ell}$ be the function value accrued at leaf node $\ell$. For any realization $\{x_{w}:w\in\rho\}$, we use ($\ell(\{x_{w}:w\in\rho\})$ to denote the (unique) leaf that is reached when subtree $\rho$ is executed under this realization: this corresponds to following the branch labeled $x_{w}$ out of each node $w\in\rho$ (starting from node $v$). We now define function $h$ that maps any realization $\{x_{w}:w\in\rho\}$ to the function value $f_{k}$ at the leaf $k=\ell(\{x_{w}:w\in\rho\})$. By our assumption on the objective function of the stochastic problem, $h$ is bounded between 00 and $f_{max}$. We now define two product distributions on nodes of $\rho$:
$\mathbf{{V}}_{1}=\langle\mathcal{D}_{v},\{\mathcal{E}_{w}:w\in\rho\setminus v\}\rangle\quad\mbox{ and }\quad\mathbf{{V}}_{2}=\langle\mathcal{E}_{v},\{\mathcal{E}_{w}:w\in\rho\setminus v\}\rangle.$
Note that
$h(\mathbf{{V}}_{1})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\quad\mbox{ and }\quad h(\mathbf{{V}}_{2})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]$
Applying Lemma 1.3 to function $h$ and product distributions $\mathbf{{V}}_{1}$ and $\mathbf{{V}}_{2}$,
$\bigg{|}\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\bigg{|}\,\,\leq\,\,f_{max}\cdot\mathtt{TV}(\mathbf{{V}}_{1},\mathbf{{V}}_{2})\,\leq\,f_{max}\cdot\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v}).$ (16)
The last inequality is by Lemma 1.2 and the fact that $\mathbf{{V}}_{1}$ and $\mathbf{{V}}_{2}$ only differ at node $v$.
We now combine the conditional expectations from (15) and (16). Using (14), the probability of event ${\cal R}_{v}$ under both distributions $\mathbf{{H}}^{v-1}$ and $\mathbf{{H}}^{v}$ is the same, which equals $Q_{v}(\sigma)$. So,
$\mid f(\sigma\mid\mathbf{{H}}^{v})-f(\sigma\mid\mathbf{{H}}^{v-1})\mid$ $=\bigg{|}\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\right]\bigg{|}$
$=Q_{v}(\sigma)\cdot\bigg{|}\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\bigg{|}$
$\leq Q_{v}(\sigma)\cdot f_{max}\cdot\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v})=f_{max}\cdot Q_{v}(\sigma)\cdot\epsilon_{v}.$
This completes the proof of (12) and Lemma 2.3.
2.3 Calculating a Stochastically Dominating Empirical Distribution
We now prove Theorem 2.1, that given i.i.d. samples from an unknown distribution $\mathcal{D}$, constructs (w.h.p.) a stochastically dominating distribution $\mathcal{E}$ with small total-variation distance.
In order to motivate this algorithm, consider the simple case that $\mathcal{D}$ is a Bernoulli distribution, i.e., $\mathcal{D}=\mathtt{Bern}(p)$ where $p$ is the (unknown) probability of realizing to value $1$. We first compute the empirical mean $\widehat{p}$ of $p$ from the $m$ i.i.d. samples. Then, we set $\mathcal{E}=\mathtt{Bern}\left(\widehat{p}+\sqrt{\frac{\log(1/\delta)}{m}}\right)$, where $\delta>0$ is the algorithm’s failure probability. It is easy to see, by Chernoff-Hoeffding bounds, that $|p-\widehat{p}|<\sqrt{\frac{\log(1/\delta)}{m}}$ and $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ with probability at least $1-\delta$. Indeed, this is the strategy used in many algorithms that use the principle of optimism under uncertainty, such as the upper-confidence-bound (UCB) algorithm [LS20]. We will prove Theorem 2.1 by giving such a construction for any distribution $\mathcal{D}$ with finite support. We note that similar constructions have been used before, e.g. [GHZ19], but their notion of approximating $\mathbf{D}$ is different from our $\mathtt{TV}$-distance condition.
Consider any distribution $\mathcal{D}$ with support $\{a_{1},\cdots,a_{k}\}$ where $a_{1}<\cdots<a_{k}$, and let $\{p_{i}\}_{i=1}^{k}$ denote the respective probabilities and $\mathsf{X}$ the corresponding r.v. That is, $\mathbf{P}(\mathsf{X}=a_{i})=p_{i}$ for all $i\in[k]$. The algorithm takes as input $m$ i.i.d. samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$. It defines the empirical distribution as $\widehat{p}_{i}:=\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[\mathsf{X}_{j}=a_{i}]$ for all $i\in[k]$. Since this empirical distribution might not stochastically dominate $\mathcal{D}$, the algorithm shifts some mass from lower to higher values in the support. However, this shift of mass needs to ensure that the resulting distribution is close to $\mathcal{D}$ in terms of total variation ($\mathtt{TV}$) distance. Algorithm 2 describes the entire procedure.
Algorithm 2 EmpStocDom
1:Input: support $S:=\{a_{1},\cdots,a_{k}\}$, samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$, confidence parameter $\delta$
2:let $\epsilon=\sqrt{\frac{\log(2k/\delta)}{2m}}$
3:for $i\in[k]$, let $\widehat{p}_{i}\leftarrow\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[\mathsf{X}_{j}=a_{i}]$
4:set $\bar{p}_{k}=\min(\widehat{p}_{k}+\epsilon,1)$
5:if $\bar{p}_{k}$ = 1 then
6:     set $\bar{p}_{i}=0$ for all $i\in[k-1]$
7:else
8:     let $y$ be the minimum index such that $\sum_{i=1}^{y}\widehat{p}_{i}\geq\epsilon$, and set
$\bar{p}_{i}=\begin{cases}\widehat{p}_{i}&\text{if }y<i\leq k-1\\ \sum_{j=1}^{y}\widehat{p}_{j}-\epsilon&\text{ if }i=y\\ 0&\text{if }i<y\end{cases}$ (17)
9:Output: distribution $\mathcal{E}$ over support $S$ with probabilities $\{\bar{p}_{i}\}_{i=1}^{k}$
We will now show that the empirical distribution $\mathcal{E}$ returned by our algorithm stochastically dominates $\mathcal{D}$ and is also a good approximation in terms of $\mathtt{TV}$ distance.
$\mathcal{E}$ is a valid distribution.
We will show that $\bar{p}_{i}\geq 0$ for all $i\in[k]$ and $\sum_{i=1}^{k}\bar{p}_{i}=1$. If $\bar{p}_{k}=1$ then it is easy to see that $\mathcal{E}$ is a valid: all other probabilities $\bar{p}_{i}=0$. If $\bar{p}_{k}<1$ then $\widehat{p}_{k}<1-\epsilon$, which implies $\sum_{i=1}^{k-1}\widehat{p}_{i}=1-\widehat{p}_{k}>\epsilon$. So, index $y\in[k-1]$, and
$\sum_{i=1}^{k}\bar{p}_{i}=\bar{p}_{y}+\sum_{i=y+1}^{k-1}\bar{p}_{i}+\bar{p}_{k}=\sum_{i=1}^{y}\widehat{p}_{i}-\epsilon+\sum_{i=y+1}^{k-1}\widehat{p}_{i}+\widehat{p}_{k}+\epsilon=1.$
Stochastic dominance and TV bound.
For any $i\in[k]$, let $q_{i}=\sum_{j\geq i}p_{j}$ and $\widehat{q}_{i}=\sum_{j\geq i}\widehat{p}_{j}$. Using Hoeffding’s bound and $\epsilon=\sqrt{\frac{\log(2k/\delta)}{2m}}$,
$\mathbf{P}\left(|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq 2e^{-2m\epsilon^{2}}\leq\frac{\delta}{k},\qquad\forall i\in[k].$
Using the union bound we have,
$\mathbf{P}\left(\exists i\in[k]\,:\,|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq\delta.$
Henceforth, we assume the “good” event that $|\widehat{q}_{i}-q_{i}|<\epsilon$ for all $i\in[k]$, which occurs with probability at least $1-\delta$. Under this event, we will show that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ and $\mathtt{TV}(\mathcal{E},\mathcal{D})<2k\epsilon$. This would complete the proof of Theorem 2.1.
Bounding TV distance between $\mathcal{D}$ and the empirical distribution.
Let $\widehat{\mathcal{E}}$ denote the distribution with probabilities $\{\widehat{p}_{i}\}_{i=1}^{k}$, which is just the empirical distribution. We will show that (under the good event),
$\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})<(k-1)\epsilon.$ (18)
Let $q_{k+1}=\widehat{q}_{k+1}=0$. Then, we have $\widehat{p}_{i}=\widehat{q}_{i}-\widehat{q}_{i+1}$ and $p_{i}=q_{i}-q_{i+1}$ for all $i\in[k]$. So,
$|\widehat{p}_{i}-p_{i}|=|\widehat{q}_{i}-q_{i}+q_{i+1}-\widehat{q}_{i+1}|\leq|\widehat{q}_{i}-q_{i}|+|q_{i+1}-\widehat{q}_{i+1}|<2\epsilon,\quad\forall i\in[k].$
Furthermore, we have $|\widehat{p}_{k}-p_{k}|<\epsilon$ and $|\widehat{p}_{1}-p_{1}|<\epsilon$, because $q_{k+1}=\widehat{q}_{k+1}=0$ and $q_{1}=\widehat{q}_{1}=1$. So,
$\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\widehat{p}_{i}-p_{i}|<(k-1)\epsilon,$
which proves (18).
Bounding TV distance between $\widehat{\mathcal{E}}$ and $\mathcal{E}$.
We consider two cases:
  • If $\bar{p}_{k}=1$ then we have $\widehat{p}_{k}\geq 1-\epsilon$ and $\bar{p}_{i}=0$ for $i\in[k-1]$, which implies:
    $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|=\frac{1}{2}\left(|1-\widehat{p}_{k}|+\sum_{i=1}^{k-1}\widehat{p}_{i}\right)\leq\epsilon.$
  • If $\bar{p}_{k}<1$ then $\bar{p}_{k}-\widehat{p}_{k}=\epsilon$ and using the definition of $\bar{p}$ from (17),
    $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|=\frac{1}{2}\left(\epsilon+|\sum_{i=1}^{y}\widehat{p}_{i}-\epsilon-\widehat{p}_{y}|+\sum_{i=1}^{y-1}\widehat{p}_{i}\right)=\epsilon,$
    where the last equality used the choice of index $y$ to conclude $|\sum_{i=1}^{y}\widehat{p}_{i}-\epsilon-\widehat{p}_{y}|=\epsilon-\sum_{i=1}^{y-1}\widehat{p}_{i}$.
Hence, in either case, we have $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})\leq\epsilon$, which combined with (18) implies
$\mathtt{TV}(\mathcal{E},\mathcal{D})\leq\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})+\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})<k\epsilon.$
Showing $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$.
Again, we consider two cases:
  • If $\bar{p}_{k}=1$ then $\mathcal{E}$ puts all its mass on the highest value $a_{k}$: so it is clear that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$.
  • If $\bar{p}_{k}<1$, consider any $v\in\mathbb{R}$. If $v\leq a_{y}$,
    $\mathbf{P}(\mathsf{Z}\geq v)=1\geq\mathbf{P}(\mathsf{X}\geq v),$
    where $\mathsf{X}$ and $\mathsf{Z}$ have distributions $\mathcal{D}$ and $\mathcal{E}$ respectively. If $v>a_{y}$, then let $i\geq y+1$ be the smallest index such that $a_{i}\geq v$. We have
    $\mathbf{P}(\mathsf{Z}\geq v)=\sum_{j=i}^{k}\bar{p}_{j}=\sum_{j=i}^{k-1}\widehat{p}_{j}+\widehat{p}_{k}+\epsilon=\widehat{q}_{i}+\epsilon\geq q_{i}=\mathbf{P}(\mathsf{X}\geq v)\,,$
    where the inequality follows due to the good event. Hence, $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$.
2.4 Framework for Down-Monotone Problems
The algorithmic framework described above assumed that the stochastic problem is up-monotone. We now describe the changes needed to handle down-monotone problems. See Definition 1.1 for the formal definition of up/down monotonicity.
The only change to Algorithm 1 is that, at each step $t$, we now need to compute distribution $\mathbf{E}^{t}$ that is stochastically dominated by the true (unknown) distribution $\mathbf{D}$. That is, we need $\mathbf{E}^{t}\preceq_{\mathsf{SD}}\mathbf{D}$ rather than $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}^{t}$. Such a distribution can be constructed in exactly the same way as Algorithm 2: we just order the values in the support in decreasing order $a_{1}>\cdots>a_{k}$ (instead of increasing order). Using the same proof as in §2.3, we obtain:
Theorem 2.4.
There is an algorithm that, given $m$ i.i.d. samples from a distribution $\mathcal{D}$ with finite support-size $k$, and parameter $\delta>0$, computes a distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$:
  • $\mathcal{E}$ is stochastically dominated by $\mathcal{D}$, and
  • the total-variation distance $\mathtt{TV}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2k/\delta)}{2m}}$.
For the analysis, we use the following analogue of the stablility lemma with down-monotonicity.
Lemma 2.5.
Consider a stochastic problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{E}\preceq_{\mathsf{SD}}\mathbf{D}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$, and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
$f(\sigma)-\alpha\cdot f(\sigma^{*})=\mathbb{E}_{\mathbf{x}\sim\mathbf{D}}\left[f(\sigma,\mathbf{x})-\alpha\cdot f(\sigma^{*},\mathbf{x})\right]\leq f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma)\cdot\epsilon_{i}\,,$
where ${Q}_{i}(\sigma)$ denotes the probability that item $i$ is probed by policy $\sigma$ under distribution $\mathbf{{D}}$.
The proof of this lemma is identical to that in §2.2 for up-monotonicity. Using Theorem 2.4 and Lemma 2.5, the proof of Theorem 1.1 (for down-monotone problems) is then identical to that for up-monotonicity (done earlier).
2.5 Framework for Maximization Problems
The algorithm/analysis so far assumed that the stochastic problem $\mathcal{P}$ has a minimization objective. We now describe the changes needed to handle a maximization problem $\mathcal{P}$. Given a product distribution $\mathbf{{U}}$, we have
$\mathtt{OPT}(\mathbf{{U}})=\max_{\tau\in\mathcal{C}}f(\tau|\mathbf{{U}}),$
where $\mathcal{C}$ is the set of all feasible policies and $f(\tau|\mathbf{{U}}):=\mathbb{E}_{\mathbf{x}\sim\mathbf{{U}}}[f(\tau,\mathbf{x})]$ is the expected value of policy $\tau$ under distribution $\mathbf{{U}}$. Note that $f(\tau)=f(\tau|\mathbf{D})$ where $\mathbf{D}$ is the true distribution. As before, we assume access to a $\alpha$-approximation algorithm $\mathtt{ALG}$ for problem $\mathcal{P}$; that is, given any problem instance (with item distributions), $\mathtt{ALG}$ finds a policy of expected value at least $\alpha$ times the optimum; here $\alpha\leq 1$. Furthermore, we assume that $\mathcal{P}$ is down-monotone (up-monotone problems can be handled similarly). We note that the algorithm is unchanged: we still use Algorithm 1. The main change is in the expression for regret, which is reflected in the modified stability lemma:
Lemma 2.6.
Consider a stochastic maximization problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
$\alpha\cdot f(\sigma^{*})-f(\sigma)=\mathbb{E}_{\mathbf{x}\sim\mathbf{D}}\left[\alpha\cdot f(\sigma^{*},\mathbf{x})-f(\sigma,\mathbf{x})\right]\leq f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma)\cdot\epsilon_{i}\,,$ (19)
where ${Q}_{i}(\sigma)$ denotes the probability that item $i$ is probed by policy $\sigma$ under distribution $\mathbf{{D}}$.
The proof of this lemma follows the same approach as for Lemma 2.3.
$\alpha\cdot f(\sigma^{*})-f(\sigma)$ $=\alpha\cdot f(\sigma^{*}|\mathbf{D})-f(\sigma|\mathbf{D})=\alpha\cdot f(\sigma^{*}|\mathbf{D})-f(\sigma|\mathbf{E})+f(\sigma|\mathbf{E})-f(\sigma|\mathbf{D})$
$\leq\alpha\cdot f(\sigma^{*}|\mathbf{D})-\alpha\cdot\mathtt{OPT}(\mathbf{E})+f(\sigma|\mathbf{E})-f(\sigma|\mathbf{D})$ (20)
$=\alpha\cdot\mathtt{OPT}(\mathbf{D})-\alpha\cdot\mathtt{OPT}(\mathbf{E})+f(\sigma|\mathbf{E})-f(\sigma|\mathbf{D})\,\,\leq\,\,f(\sigma|\mathbf{E})-f(\sigma|\mathbf{D})$ (21)
$\leq f_{max}\cdot\sum_{i=1}^{n}\epsilon_{i}\cdot Q_{i}(\sigma).$ (22)
Inequality (20) uses the fact that $\sigma$ is an $\alpha$-approximate policy to the instance with distribution $\mathbf{E}$. In (21), the equality uses the fact that $\sigma^{*}$ is an optimal policy for the instance with distribution $\mathbf{D}$, and the inequality is by down-monotonicity (Definition 1.1) and $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$. Finally, (22) follows from the upper bound on $\bigg{|}f(\sigma|\mathbf{D})-f(\sigma|\mathbf{E})\bigg{|}$ from (13).
Lemma 2.6 then implies Theorem 1.1 for maximization, down-monotone problems: the proof is identical to that for minimization problems (see §2.1). Finally, the same guarantee holds when $\mathcal{P}$ is up-monotone (per the discussion in §2.4).
2.6 Framework for Continuous Distributions
The algorithmic framework described above assumed that the random variables are discrete. Here, we show how to handle stochastic problems with arbitrary (continuous or discrete) r.v.s as long as the policies are based on a finite number of thresholds. We first define a natural discretization of continuous r.v.s.
Definition 2.1 (Thresholded r.v.).
For any $[0,U]$-bounded r.v. $\mathsf{X}$ (possibly continuous) and list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of increasing threshold values, let $\theta(\mathsf{X},\mathbf{{b}})$ be the r.v. such that
$\theta(\mathsf{X},\mathbf{{b}})=i\text{ if }b_{i-1}\leq\mathsf{X}<b_{i},\qquad\forall i\in[k],$
where we use $b_{0}:=0$ and $b_{k}:=U$. This corresponds to thresholding $\mathsf{X}$ according to $\mathbf{{b}}$.
Note that $\theta(\mathsf{X},\mathbf{{b}})$ is a discrete random variable with support size $k$, even if $\mathsf{X}$ is continuous. We will assume that the $\alpha$-approximation algorithm $\mathtt{ALG}$ for the stochastic problem $\mathcal{P}$ always finds a $k$-threshold policy, defined as follows. Such a policy is given by a decision tree $\sigma$, where each node is labeled by an item $j$ to probe and list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of thresholds, such that branch $i\in\{1,\cdots k\}$ is followed when $\mathsf{X}_{j}\in[b_{i-1},b_{i})$. This limits the number of branches out of each node to at most $k$, whereas an arbitrary policy may have an unbounded number of branches (corresponding to each possible outcome). We emphasize that the thresholds at different nodes may be completely different. This restriction allows us to analyze policies involving continuous r.v.s in the same manner as before (for discrete r.v.s). In particular, all other definitions related to the policy (terminal/leaf nodes, policy execution, cost etc) remain the same as for discrete r.v.s. Moreover, for problems like Prophet inequality and Pandora’s Box, there are optimal $1$-threshold policies: so this restriction is without loss of generality for these problems.
Below, we focus on problems involving continuous r.v.s. that are up-monotone and have a minimization objective. We note that this approach also works directly for down-monotone problems or maximization objectives. We need to modify the method to compute stochastically dominating distributions and the stability lemma: these changes are described below.
Theorem 2.7.
There is an algorithm that, given $m$ i.i.d. samples from an any bounded distribution $\mathcal{D}$ and parameter $\delta>0$, computes a (discrete) distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$:
  • $\mathcal{E}$ is stochastically dominates $\mathcal{D}$.
  • For any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of $k-1$ thresholds, we have
    $\mathtt{TV}(\theta(\mathsf{X},\mathbf{{b}}),\theta(\mathsf{Y},\mathbf{{b}}))<k\sqrt{\frac{\log(2(k+m)/\delta)}{2m}},\,\,$
    where $\mathsf{X}$ (resp. $\mathsf{Y}$) is distributed according to $\mathcal{D}$ (resp. $\mathcal{E}$).
Proof Outline.
The algorithm is a natural extension of that for the discrete case, and is described in Algorithm 3. The key idea is to put an $\epsilon$ mass on the highest value $U$ and remove an equal mass from the lowest sampled values.
Algorithm 3 StocDomCont
1:Input: samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$, range $[0,U)$, confidence parameter $\delta$
2:let $\epsilon=\min\left\{\sqrt{\frac{\log(2(k+m)/\delta)}{2m}},1\right\}$
3:let $a_{1}\leq\cdots\leq a_{m}$ denote the sampled values, and $a_{m+1}:=U$
4:let $\widehat{p}_{i}=\frac{1}{m}$ for $i\in[m]$ and $\widehat{p}_{m+1}=0$ denote the sample distribution
5:let $y$ be the minimum index such that $\sum_{i=1}^{y}\widehat{p}_{i}\geq\epsilon$, and set
$\bar{p}_{i}=\begin{cases}\widehat{p}_{i}+\epsilon&\text{if }i=m+1\\ \widehat{p}_{i}&\text{if }y<i\leq m\\ \sum_{j=1}^{y}\widehat{p}_{j}-\epsilon&\text{if }i=y\\ 0&\text{if }i<y\end{cases}$ (23)
6:Output: distribution $\mathcal{E}$ with support $\{a_{i}\}_{i=1}^{m+1}$ and probabilities $\{\bar{p}_{i}\}_{i=1}^{m+1}$
It is easy to see that $\mathcal{E}$ is a valid distribution: note that $\bar{p}\geq 0$ and $\sum_{i=1}^{m+1}\bar{p}_{i}=1$. We are now ready to prove the two properties. Consider any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of thresholds for the second property. Let $c_{1}\leq\cdots\leq c_{r}$ be the values in $\{a_{i}\}_{i=1}^{m+1}\cup\{b_{j}\}_{j=1}^{k-1}$ in increasing order. Note that the $c_{i}$s contain both the sampled values and the thresholds, and $r\leq m+k$. Moreover, the last value $c_{r}=a_{m+1}=U$. We extend the probabilities in $\widehat{p}$ and $\bar{p}$ to all values $c_{1}\leq\cdots\leq c_{r}$ (just set the probability to zero for non-sampled values). To keep notation simple, we will still refer to these “extended” distributions as $\widehat{p}$ and $\bar{p}$.
We again use the empirical distribution as an intermediate quantity: let $\widehat{\mathcal{E}}$ denote the distribution with support $\{c_{i}\}_{i=1}^{r}$ and probabilities $\{\widehat{p}_{i}\}_{i=1}^{r}$; let $\mathsf{Z}$ denote the corresponding r.v. For any $i\in[r]$, let $q_{i}=\mathbf{P}(\mathsf{X}\geq c_{i})$ and $\widehat{q}_{i}=\mathbf{P}(\mathsf{Z}\geq c_{i})$. Using Hoeffding’s bound and union bound (as before), and the value $\epsilon=\sqrt{\frac{\log(2(m+r)/\delta)}{2m}}$,
$\mathbf{P}\left(\exists i\in[r]\,:\,|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq\delta.$
Henceforth, we assume the “good” event that $|\widehat{q}_{i}-q_{i}|<\epsilon$ for all $i\in[r]$, which occurs with probability at least $1-\delta$. We will prove the claimed properties under this event, which would complete the proof of Theorem 2.7.
Stochastic Dominance.
Recall that $y\in[m]$ is the smallest index among the sampled values where the “prefix” probability until $y$ is at least $\epsilon$. As we are working with the expanded set of $c_{i}$ values, for the analysis, we redefine $y\in[r]$ to be the smallest index with $\sum_{i=1}^{y}\widehat{p}_{i}\geq\epsilon$. Now, consider any $v\in\mathbb{R}$. If $v\leq c_{y}$,
$\mathbf{P}(\mathsf{Y}\geq v)=1\geq\mathbf{P}(\mathsf{X}\geq v),$
where $\mathsf{Y}$ has distribution $\mathcal{E}$. If $v>c_{y}$, then let $i\geq y+1$ be the smallest index such that $c_{i}\geq v$. We have
$\mathbf{P}(Y\geq v)=\sum_{j=i}^{r}\bar{p}_{j}=\sum_{j=i}^{r-1}\widehat{p}_{j}+\widehat{p}_{r}+\epsilon=\widehat{q}_{i}+\epsilon\geq q_{i}=\mathbf{P}(\mathsf{X}\geq v)\,,$
where the inequality follows due to the good event. Hence, $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$, which proves the first property.
Bounding TV distance.
For the second property, recall that $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$. We also use $b_{0}=0$ and $b_{k}=U$. Let $\mathsf{X}\sim\mathcal{D}$, $\mathsf{Z}\sim\widehat{\mathcal{E}}$, $\mathsf{Y}\sim\mathcal{E}$, and $\theta_{D}=\theta(\mathsf{X},\mathbf{{b}})$, $\theta_{F}=\theta(\mathsf{Z},\mathbf{{b}})$, $\theta_{E}=\theta(\mathsf{Y},\mathbf{{b}})$ be the corresponding thresholded r.v.s. We also define
$q^{*}_{j}=\mathbf{P}(\mathsf{X}\geq b_{j-1})=\mathbf{P}(\theta_{D}\geq j),\quad\forall j\in[k].$
The probabilities $\widehat{q}^{*}$ are defined similarly for r.v. $\mathsf{Z}$. By the good event, we have $|\widehat{q}^{*}_{j}-q^{*}_{j}|<\epsilon$ for all $j\in[k]$ as all the thresholds are included in the $c_{i}$ values. Now, exactly as for the discrete case, we obtain:
$\mathtt{TV}(\theta_{D},\theta_{F})<(k-1)\epsilon.$
Now, we bound the distance between $\theta_{E}$ and $\theta_{F}$ using the fact that r.v.s $\mathsf{Y}$ and $\mathsf{Z}$ are discrete:
$\mathtt{TV}(\theta_{F},\theta_{E})\leq\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{r}|\bar{p}_{i}-\widehat{p}_{i}|<\epsilon,$
where the last inequality uses the choice of index $y$ (as done in the discrete case). Combining the above two inequalities, we get $\mathtt{TV}(\theta_{D},\theta_{F})<k\epsilon$, which proves the second property. ∎
Next, we state the modified version of the stability lemma.
Lemma 2.8.
Consider a stochastic problem that is up-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions, where for each $i\in[n]$, distributions $\mathcal{D}_{i}$ and $\mathcal{E}_{i}$ satisfy:
  • $\mathcal{E}_{i}$ stochastically dominates $\mathcal{D}_{i}$.
  • For any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of $k-1$ thresholds, we have
    $\mathtt{TV}(\theta(\mathsf{X},\mathbf{{b}}),\theta(\mathsf{Y},\mathbf{{b}}))<\epsilon_{i},$
    where $\mathsf{X}$ (resp. $\mathsf{Y}$) is distributed according to $\mathcal{D}_{i}$ (resp. $\mathcal{E}_{i}$).
If $\sigma=\mathtt{ALG}(\mathbf{E})$ is a $k$-threshold policies and $\sigma^{*}$ is the optimal policy under $\mathbf{D}$, then:
$f(\sigma)-\alpha\cdot f(\sigma^{*})=\mathbb{E}_{\mathbf{x}\sim\mathbf{D}}\left[f(\sigma,\mathbf{x})-\alpha\cdot f(\sigma^{*},\mathbf{x})\right]\leq f_{\max}\sum_{i=1}^{n}{Q}_{i}(\sigma)\cdot\epsilon_{i}\,,$
where ${Q}_{i}(\sigma)$ denotes the probability that item $i$ is probed by policy $\sigma$ under distribution $\mathbf{{D}}$.
The proof of this lemma is almost identical to that of Lemma 2.3 (see §2.2), so we just explain the changes. Recall that policy $\sigma$ corresponds to a decision tree. Using the additional fact that $\sigma$ is a $k$-threshold policy, it follows that the execution of $\sigma$ only depends on the “thresholded” r.v.s $\{\theta_{\sigma(v)}:v\in[N]\}$ where $\theta_{\sigma(v)}$ is the result of applying the thresholds (at most $k$) at node $v$ on the continuous r.v. $\mathsf{X}_{\sigma(v)}$. Crucially, these $\theta$ r.v.s are discrete and independent: so we can directly use the proof of Lemma 2.3 (where we replace the $X_{v}$ r.v.s by $\theta_{\sigma(v)}$).
Using Theorem 2.7, Lemma 2.8, and the same steps as for proving Theorem 1.1, we obtain:
Theorem 2.9.
Suppose that stochastic problem $\mathcal{P}$ has an $\alpha$-approximation algorithm via $k$-threshold policies, and $\mathcal{P}$ is either up-monotone or down-monotone. Then, there is a polynomial-time semi-bandit learning algorithm for $\mathcal{P}$ with $\alpha$-regret $O(knf_{max}\sqrt{T\log(knT)})$. Here, $n$ is the number of items, $f_{max}$ is the maximal value of the objective function and $T$ is the number of periods.
3 Applications
In this section, we show that several stochastic optimization problems are covered by our framework, resulting in $\sqrt{T\log T}$ regret online learning algorithm for all these problems.
Series Testing
We start with a simple problem: there are $n$ components, where each component $i$ is “working” independently with some known probability $p_{i}$. To determine if any component $i\in[n]$ is working, we need to perform a test, which costs $c_{i}$. All $n$ components must be working for the system to be functional. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Note that testing stops once a failed component is found: so we do not observe all the outcomes and only have semi-bandit feedback. It is easy to show that this problem is up-monotone: this is a (simple) special case of Lemma 3.4 below. It is well-known that the natural greedy policy achieves the optimal cost [But72]. So, using Theorem 1.1 with $k=2$ (all r.v.s are binary), we obtain a polynomial time online learning algorithm for series testing having $1$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
Prophet Inequality
The Prophet Inequality [KS77, SC84] is a fundamental problem in optimal stopping, which has also been used extensively in algorithmic game theory. The input consists of $n$ rewards which arrive in a given fixed sequence, say $\mathsf{X}_{1},\cdots\mathsf{X}_{n}$. Each reward $\mathsf{X}_{i}$ is drawn independently from a known distribution $\mathcal{D}_{i}$. We are interested in online policies, that upon observing each reward, selects or discards it immediately. The policy can select at most one reward, and it terminates right after making a selection (without observing any future reward). Note that we have semi-bandit feedback because only some of the rewards are observed in any policy execution. The goal is to maximize the expected selected reward. The classical results obtain a $\frac{1}{2}$-approximate policy relative to the “clairvoyant” optimal value $\mathbb{E}[\max_{i=1}^{n}\mathsf{X}_{i}]$; there are also instances where no policy can achieve a better approximation to this benchmark. Here, we will compare to a more realistic non-clairvoyant benchmark: the optimal policy which is also constrained to make selection decisions in the given order (same as an algorithm). It is known that there is an optimal threshold-based policy: given thresholds $\{\tau_{i}\}_{i=1}^{n}$, the policy selects $i$ if and only if $\mathsf{X}_{i}>\tau_{i}$. This is a $1$-threshold policy, as defined in §2.6. Moreover, the prophet inequality problem is strongly monotone (see Lemma 28 in [GHTZ21]), which implies that it is down-monotone. Using Theorem 2.9, we get a polynomial time online learning algorithm for the the prophet inequality problem with unknown distributions that has $1$-regret $O(nU\sqrt{T\log T})$ where all r.v.s are $[0,U]$ bounded. This improves upon the $O(n^{3}U\sqrt{T}\log T)$ bound from [GKSW24], although the previous result holds in the stronger bandit-feedback model. We note that there are other learning-based results [AKW14, RWW20] based on limited number of samples, that imply $\frac{1}{2}$-regret algorithms by comparing to the clairvoyant benchmark. Note that our guarantee and that of [GKSW24] are stronger because they do not incur any multiplicative approximation factor.
Combinatorial Prophet Inequalities
The basic prophet inequality concept has also been extended to settings where there is some combinatorial feasibility constraint on the selected items. Here, we have $n$ items with reward $\mathsf{X}_{i}\sim\mathcal{D}_{i}$ for each $i\in[n]$. In addition, there is a downward-closed [^4] Set family ${\cal F}$ is downward-closed if, for any $S\in{\cal F}$ we have $S^{\prime}\in{\cal F}$ for every $S^{\prime}\subseteq S$. set family ${\cal F}\subseteq 2^{[n]}$ that represents a generic feasibility constraint; the selected subset must be in ${\cal F}$. The $n$ items arrive in a given fixed sequence. When item $i$ arrives, if $S\cup\{i\}\not\in{\cal F}$ where $S$ is the set of previously selected items then item $i$ is not considered for selection (and we do not observe $\mathsf{X}_{i}$); otherwise, the policy observes the value of $\mathsf{X}_{i}$ and selects/discards item $i$. Again, note that we only have semi-bandit feedback because only a subset of items is observed by the policy. The performance of an online policy is compared to the clairvoyant optimum $\mathtt{OPT}^{*}=\mathbb{E}\left[\max_{S\in{\cal F}}\sum_{i\in S}\mathsf{X}_{i}\right]$. Many specific problems can be modeled in this manner:
  • When =${\cal F}=\{S:|S|\leq 1\}$, we get the classic prophet inequality, which has a $\frac{1}{2}$ approximation.
  • When =${\cal F}=\{S:|S|\leq k\}$, there is a $1-\frac{1}{\sqrt{k+3}}$ approximation [Ala14, JMZ22].
  • When ${\cal F}$ corresponds to independent sets in a matroid, we obtain the matroid prophet inequality where again a $\frac{1}{2}$ approximation is known [KW19].
  • When ${\cal F}$ is given by the intersection of $p$ matroids, a $\frac{1}{e(p+1)}$ approximation is known [FSZ21].
  • When ${\cal F}$ is given by matchings in a graph, a $0.337$ approximation is known [EFGT22].
The online policies in all these results are $1$-threshold policies (as in §2.6). For the monotone property, note that we are comparing to the clairvoyant optimum $\mathtt{OPT}^{*}$ (not the optimal policy). So, it suffices to prove down-monotonicity for $\mathtt{OPT}^{*}$, which is immediate by stochastic dominance. Therefore, using Theorem 2.9, we obtain $\alpha$-regret $O(nU\sqrt{T\log T})$ for all the above combinatorial prophet inequalities where $\alpha$ is the best (offline) approximation ratio; we assume that the r.v.s are $[0,U]$ bounded.
Some approximate regret bounds can also be obtained from previous work on single-sample prophet inequalitites [AKW14, RWW20, ${}^{+}$22]. While the regret bounds via this approach are better (just $O(n)$), they need to compare to approximation ratios that are often much worse than the usual (known distribution) setting. For example, no constant-factor approximation is known for single-sample matroid prophet inequality, whereas our results imply $\frac{1}{2}$-regret of $\sqrt{T\log T}$ for general matroids, matching the approximation ratio from [KW19].
Pandora’s Box.
In this problem [Wei79], we are given distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ such that r.v. $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. The realization of $\mathsf{X}_{i}$ can be ascertained by paying a known inspection cost $c_{i}$. Now, the goal is to find a policy to (adaptively) inspect a subset $S\subseteq[n]$ of the random variables to maximize $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. Note that any policy only inspects a subset of items and we only receive feedback from these items, which corresponds to semi-bandit feedback. [Wei79] obtained an optimal policy based on the “reservation value” for each item and probing items according to this value until the reward for an item exceeds all remaining reservation values. The reservation value $r_{i}$ for an item $i$ is such that $\mathbb{E}[(\mathsf{X}_{i}-r_{i})_{+}]=c_{i}$. We note that this optimal policy is $1$-threshold, according to the definition in §2.6.
For the online setting, we assume that the r.v.s $\mathsf{X}_{i}$ are $[0,U]$ bounded; the distributions may be discrete or continuous. It was shown in [GHTZ21] (see Lemma 31 in that paper) that the Pandora’s box problem is strongly monotone, which implies that it is down-monotone. Combined with Theorem 2.9, we get a polynomial time online learning algorithm for the Pandora’s box problem with unknown distributions that has $1$-regret $O(n(C+U)\sqrt{T\log T})$ where $C=\sum_{i\in[n]}c_{i}$ is the total cost. Our regret bound improves upon the $O(n^{4.5}(C+U)\sqrt{T}\log T)$ bound from [GKSW24], although the previous result holds in the stronger bandit-feedback model.
Variants of Pandora’s Box.
Our framework also applies to more general versions of Pandora’s problem that have been studied in prior work. We mention two such variants here.
In Pandora’s box with order constraints, in addition to the $n$ r.v.s, there are precedence constraints that enforce that any r.v. $\mathsf{X}_{i}$ may be selected only after all its predecessors have been selected. Although the original policy [Wei79] does not apply to this extension, [BFLL23] obtained a different optimal policy when the precedence constraints form a directed tree: this policy is also a 1-threshold policy. Hence, Theorem 2.9 implies an online learning algorithm for Pandora’s box with tree order constraints having $1$-regret $O(n(C+U)\sqrt{T\log T})$, where $C=\sum_{i\in[n]}c_{i}$ is the total cost and $U$ is the bound on the r.v.s.
In the matroid Pandora’s box problem [KWW16, Sin18], in addition to the $n$ r.v.s $\{\mathsf{X}_{i}\}_{i=1}^{n}$, there is a matroid $\cal M$ with groundset $[n]$. The goal is to inspect a subset $S\subseteq[n]$ of r.v.s and select a subset $B\subseteq S$ such that $B$ is independent in matroid $\cal M$. The objective is to maximize $\mathbb{E}\left[\sum_{j\in B}\mathsf{X}_{j}-\sum_{i\in S}c_{i}\right]$, the difference between the total selected value and inspection cost. We recover the original Pandora’s box problem when $\cal M$ is a rank-1 uniform matroid. There is an optimal 1-threshold policy known for this variant [Sin18]. This policy is also non-adaptive. So, by Theorem 2.9, we obtain an online learning algorithm for matroid Pandora’s box having $1$-regret $O(n(C+U)\sqrt{T\log T})$, where $C=\sum_{i\in[n]}c_{i}$ is the total cost and $U$ is the bound on the r.v.s.
Stochastic Knapsack
This is a classic problem in stochastic optimization, which was introduced in [DGV08] and has been studied extensively [BGK11, GKMR11, Ma18]. There are $n$ items with deterministic rewards $\{r_{i}\}_{i=1}^{n}$ and random costs $\{C_{i}\sim\mathcal{D}_{i}\}_{i=1}^{n}$. The realized cost $C_{i}$ of item $i$ is only known after selecting it. Given a knapsack budget $B$, a policy selects items sequentially until the total cost exceeds $B$. The objective is to maximize the expected total reward from items that fit in the knapsack. (If there is an item that overflows the budget then it does not contribute to the objective.) Note that only some subset of items is selected by a policy, and we only observe those costs as feedback. In Lemma 3.1 below, we show that this problem is down-monotone (assuming that the costs are discrete r.v.s).
There is a non-adaptive $\frac{1}{4}$-approximation algorithm for stochastic knapsack [DGV08], which probes items in a fixed sequence until the budget is exhausted; so it is a 1-threshold policy. Theorem 2.9 then implies a polynomial time online learning algorithm for stochastic knapsack having $\frac{1}{4}$-regret of $O(nR\sqrt{T\log T})$ where $R=\sum_{i=1}^{n}r_{i}$ is the total reward. There is also a better adaptive algorithm for stochastic knapsack, which is a $(\frac{1}{2}-\epsilon)$ approximation (for any constant $\epsilon>0$) [Ma18]. Theorem 1.1 then implies an online learning algorithm with $(\frac{1}{2}-\epsilon)$-regret of $O(nkR\sqrt{T\log(kT)})$ where $k$ is the maximum support size.
Our results also apply to the more general stochastic orienteering problem, where items are located at vertices in a metric space and we want to find a path with budget $B$ on the total distance (from the edges in the path) plus cost (of the visited items). There is a non-adaptive $\Omega(\frac{1}{\log\log B})$ approximation algorithm for this problem [GKNR15], which is a 1-threshold policy. Here, we obtain $\Omega(\frac{1}{\log\log B})$-regret of $O(nR\sqrt{T\log T})$.
Lemma 3.1.
The stochastic (maximum) knapsack problem is up-monotone.
Proof.
We use the following result on $b$-transportations in the proof of this lemma.
$b$-Transportation. Given a bipartite graph $G=(V,E)$, and a vector $b\in\mathbb{R}^{V}_{+}$, a $b$-transportation is a vector $f\in\mathbb{R}^{E}_{+}$ with $f(\delta(v))=b_{v}$ for each $v\in V$. Here $f(\delta(v))$ denotes the sum of $f$-values of edges incident on $v$. The following theorem (due to [Rad43], see also Theorem 21.11 in [${}^{+}$03]) characterizes a $b$-transportation.
Theorem 3.2.
Let $G=(V,E)$ be a bipartite graph and let $b\in\mathbb{R}^{V}_{+}$. Then there exists a $b$-transportation if and only if $b(C)\geq\frac{1}{2}\cdot b(V)$ for every vertex cover $C$.
Consider an arbitrary instance $\mathcal{I}$ and product probability distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ where $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$. To prove the lemma, we need the following intermediate result.
Coupling distributions under stochastic dominance. Consider r.v.s $\mathsf{X}$ and $\mathsf{Y}$ with finite support $[k]=\{1,\ldots,k\}$, and distributed according to $\mathcal{D}$ and $\mathcal{E}$ respectively. Furthermore, suppose that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$. Then, we will show that there exist values $\left\{f_{\ell_{1},\ell_{2}}:k\geq\ell_{1}\geq\ell_{2}\geq 1\right\}$ such that:
$\sum_{\ell_{2}:\ell_{2}\leq\ell_{1}}f_{\ell_{1},\ell_{2}}=\mathbf{P}_{\mathcal{E}}(\mathsf{Y}=\ell_{1}),\,\forall\ell_{1}\in[k]\quad\mbox{ and }\quad\sum_{\ell_{1}:\ell_{1}\geq\ell_{2}}f_{\ell_{1},\ell_{2}}=\mathbf{P}_{\mathcal{D}}(\mathsf{X}=\ell_{2}),\,\forall\ell_{2}\in[k].$
The $\{f_{\ell_{1},\ell_{2}}\}$ values allow us to couple the realizations of r.v.s $\mathsf{X}\sim\mathcal{D}$ and $\mathsf{Y}\sim\mathcal{E}$. We can obtain these values by constructing a $b$-transportation instance on an appropriately defined bipartite graph: consider the bipartite graph $G=(U_{E}\cup U_{D},E)$ where the vertex sets represent the support of the random variable under distributions $\mathcal{E}$ and $\mathcal{D}$ respectively. Specifically, let $U_{E}=\{e_{1},\ldots,e_{k}\}$ and $U_{D}=\{d_{1},\ldots,d_{k}\}$, and $E=\{(e_{\ell_{1}},d_{\ell_{2}}):\ell_{1}\geq\ell_{2}\}$. Consider a $b$-value of $\mathbf{P}_{\mathsf{Y}\leftarrow\mathcal{E}}(\mathsf{Y}=\ell)$ on $e_{\ell}$, and a $b$-value of $\mathbf{P}_{\mathsf{X}\leftarrow\mathcal{D}}(\mathsf{X}=\ell)$ on $d_{\ell}$ for all $\ell=1,\ldots,k$. To apply Theorem 3.2, it suffices to show that for any vertex cover $C$ of $G$, we have $b(C)\leq\frac{1}{2}\cdot b(V)=1$ (since the $b$-values correspond to probabilities). Let $C$ be a fixed but arbitrary vertex cover of $G$, and observe that if $e_{\ell}\notin C$, then $C$ must contain $\{d_{\widehat{\ell}}:\widehat{\ell}\leq\ell\}$ in order to cover all edges incident on $e_{\ell}$. Consequently, any minimal vertex cover must have the form $\{e_{k},e_{k-1},\ldots,e_{\overline{\ell}},d_{\overline{\ell}-1},\ldots d_{1}\}$ where ¯$\overline{\ell}\in\{1,\ldots,k+1\}$. Let $\overline{C}\subseteq C$ be a minimal vertex cover, which must have the form described above; that is, $\overline{C}=\{e_{k},e_{k-1},\ldots,e_{\overline{\ell}},d_{\overline{\ell}-1},\ldots d_{1}\}$ for some ¯$\overline{\ell}\in\{1,\ldots,k+1\}$. Then,
$b(C)\geq b(\overline{C})=\sum_{\ell_{1}\geq\overline{\ell}}b(e_{\ell_{1}})+\sum_{\overline{\ell}>\ell_{2}\geq 1}b(d_{\ell_{2}})\geq\sum_{\ell_{1}\geq\overline{\ell}}b(d_{\ell_{1}})+\sum_{\overline{\ell}>\ell_{2}\geq 1}b(d_{\ell_{2}})=1.$
The last inequality follows from stochastic dominance $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ because:
$\sum_{\ell_{1}\geq\overline{\ell}}b(e_{\ell_{1}})=\mathbf{P}_{\mathsf{Y}\leftarrow\mathcal{E}}(\mathsf{Y}\geq\overline{\ell})\geq\mathbf{P}_{\mathsf{X}\leftarrow\mathcal{D}}(\mathsf{X}\geq\overline{\ell})=\sum_{\ell_{1}\geq\overline{\ell}}b(d_{\ell_{1}}).$
Thus, the condition of Theorem 3.2 holds, and so there exists a $b$-transportation $f:E\to\mathbf{R}$ for $G$. See Figure 1 for an example.
Figure 1: Suppose r.v.s $\mathsf{X}$ and $\mathsf{Y}$ have support $\{1,\ldots,5\}$. Let $\mathsf{X}\sim\mathcal{D}$ where $\mathcal{D}=\{1:0.2,2:0.1,3:0.1,4:0.3,5:0.3\}$. Here, each entry denotes the probability of the r.v. taking on the corresponding value; that is, $\mathbf{P}(\mathsf{X}=2)=0.1$. Similarly, let $\mathsf{Y}\sim\mathcal{E}$ where =$\mathcal{E}=\{1:0,2:0.1,3:0.3,4:0.2,5:0.4\}$. It is easy to check that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$. The figure depicts the corresponding $b$-transportation instance, and values that couple distributions $\mathcal{D}$ and $\mathcal{E}$ (as per Lemma 3.1).
We are now ready to complete the proof. For each item $i\in[n]$, let $f_{\ell_{1},\ell_{2}}^{i}$ denote coupled values for distributions $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$, constructed as per the above procedure. Let $\sigma_{\mathbf{{E}}}$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{E}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ denote its reward. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{D}}$, by closely following $\sigma_{\mathbf{{E}}}$. When policy $\sigma_{\mathbf{{E}}}$ probes some item $i$, policy $\widehat{\sigma}$ does the same. Now, suppose that $\widehat{\sigma}$ observes $C_{i}=\ell_{2}\in[k]$. Then, for each $\ell_{1}\geq\ell_{2}$, with probability $\frac{f_{\ell_{1},\ell_{2}}^{i}}{\mathbf{P}_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}$, policy $\widehat{\sigma}$ continues to follow policy $\sigma_{\mathbf{{E}}}$’s branch under the outcome $\ell_{1}$ (i.e., it wastes $\ell_{1}-\ell_{2}$ space in the knapsack). Note that the total probability $\sum_{\ell_{1}\geq\ell_{2}}\frac{f_{\ell_{1},\ell_{2}}^{i}}{\mathbf{P}_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}=1$ by definition of the $b$-transportation. Moreover, the probability of reaching any node $s$ in policy $\sigma_{\mathbf{{E}}}$ (under $\mathbf{E}$) equals the probability of reaching the same node $s$ in policy $\widehat{\sigma}$ (under $\mathbf{D}$). Thus, the expected rewards of the two policies are the same, $\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{D}}}[f(\widehat{\sigma},\mathbf{{X}})]=\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{E}}},\mathbf{{X}})]$. We note that $\widehat{\sigma}$ is a randomized policy; however, as noted in §1.4, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ as needed. ∎
Stochastic Matching and Probing
In the stochastic matching problem [${}^{+}$, ${}^{+}$12], there is an undirected graph $G=(V,E)$ with edge-weights $\{w_{e}\}_{e\in E}$, edge-probabilities $\{p_{e}\}_{e\in E}$, and vertex bounds $\{t_{v}\}_{v\in V}$. Each edge $e$ is active independently with probability $e$. However, the status (active/inactive) of an edge can only be determined by probing it. There is also a constraint on the set of probed edges: for any vertex $v$, the number of probed edges incident to $v$ must be at most $t_{v}$. A solution/policy needs to to probe a subset of edges and select a matching $M$ consisting of active edges. The objective is to maximize the expected weight of $M$. Finally, there is a “query commit” requirement that any probed edge which is active must be included in the selected matching $M$. Observe that in any policy execution, we only see the status of some subset of edges, which corresponds to semi-bandit feedback. This problem has been extensively studied, see e.g. [${}^{+}$, ${}^{+}$12, Ada11, BGMS21]. The current best approximation ratio is $0.5$ for the unweighted case [Ada11] and $0.382$ for the weighted case  [BGMS21]. We show in Lemma 3.3 below that the stochastic matching problem is down-monotone. So, Theorem 1.1 with $k=2$ (as all r.v.s are binary) implies an online learning algorithm for stochastic matching having $0.382$-regret of $O(nW\sqrt{T\log T})$ where $W=\sum_{e\in E}w_{e}$ is the total weight.
In fact, our result applies to the much more general stochastic probing problem, as defined in [GN13]. Here, we have a set $E$ of stochastic items with weights $\{w_{e}\}_{e\in E}$ and probabilities $\{p_{e}\}_{e\in E}$. Each item is active independently with probability $e$, and this status can only be determined by probing $e$. We now have two downward-closed constraints: an inner constraint ${\cal F}_{in}$ and an outer constraint ${\cal F}_{out}$. We want to probe a set $Q$ of items subject to the outer constraint (i.e. $Q\in{\cal F}_{out}$) and select a subset $S\subseteq Q$ of active (probed) items satisfying the inner constraint (i.e. $S\in{\cal F}_{in}$). The objective is to maximize the expected weight of the selected items $S$. We again have the query-commit requirement that any active probed item must be selected into the solution $S$. When both inner/outer constraints are $k_{in}$ and $k_{out}$ systems (see [GN13] for the definition of $k$-systems), there is a $\frac{1}{k_{in}+k_{out}}$ approximation for the unweighted case and an $\Omega(\frac{1}{(k_{in}+k_{out})^{2}})$ approximation for the weighted case [GN13]. When the $k$-systems are intersections of matroids, [ASW16] gave an improved adaptive algorithm with approximation ratio $\frac{1}{k_{in}+k_{out}}$, even for the weighted case.
Lemma 3.3.
The stochastic probing problem is down-monotone.
Proof.
Consider an arbitrary instance $\mathcal{I}$ and product probability distributions $\mathbf{{D}}=\{p_{e}\}_{e\in E}$ and $\mathbf{{E}}=\{\bar{p}_{e}\}_{e\in E}$ where $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$. As the r.v.s are binary in this problem, the stochastic dominance just means $p_{e}\leq\bar{p}_{e}$ for all $e\in E$.
Let $\sigma$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{D}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ denote its expected weight. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random $0-1$ realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{E}}$, by closely following $\sigma$. When policy $\sigma$ probes item $e$ at any node, policy $\widehat{\sigma}$ does the following:
  • probe $e$ with probability $p_{e}/\bar{p}_{e}$, and follow the active/inactive branch based on its outcome.
  • skip directly to the inactive branch with probability $1-p_{e}/\bar{p}_{e}$.
Note that conditioned on reaching such a node, the probability that $\widehat{\sigma}$ goes to the active branch is $\frac{p_{e}}{\bar{p}_{e}}\cdot\bar{p}_{e}=p_{e}$; so the probability that $\widehat{\sigma}$ goes to the inactive branch is $1-p_{e}$. Therefore, the probability of reaching any node $v$ in policy $\widehat{\sigma}$ (under distribution $\mathbf{E}$) equals the probability of reaching node $v$ in policy ${\sigma}$ (under distribution $\mathbf{D}$). Thus the expected values of these two policies are the same, $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\widehat{\sigma},\mathbf{{x}})]=\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{D}}}[f(\sigma,\mathbf{{x}})]$. We note that $\widehat{\sigma}$ also satisfies the query-commit requirement because $\sigma$ does (whenever $e$ is probed and found active, it will be included in the solution). Finally, observe that $\widehat{\sigma}$ is a randomized policy; however, as noted in §1.4, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ as needed. ∎
Stochastic Covering Knapsack
Here, we have $n$ items with deterministic costs $\{c_{i}\}_{i\in[n]}$ and (independent) random rewards $\{\mathsf{R}_{i}\}_{i\in[n]}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{R}_{i}$; we assume discrete distributions here. Given a target $Q$, we need to select items until the total observed reward is at least $Q$ (or we run out of items). The objective is to minimize the expected cost. Note that we only observe items until the target is achieved, which corresponds to semi-bandit feedback. There is an adaptive $3$-approximation algorithm for this problem [DHK16]. By a proof identical to that of Lemma 3.1, we get:
Lemma 3.4.
The stochastic covering knapsack problem is up-monotone.
So, using Theorem 1.1, we obtain a polynomial time online learning algorithm for stochastic min-knapsack having $3$-regret $O(nkC\sqrt{T\log(kT)})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost and $k$ is the maximum support size.
Stochastic $k$-TSP
The stochastic $k$-TSP problem is the stochastic variant of the classical $k$-TSP problem [Gar05]. An instance of the problem comprises a metric $(V,d)$ with a root vertex $r\in V$. There are random rewards $\{\mathsf{R}_{v}\}_{v\in V}$ associated with the vertices. All reward distributions are given as input, but the realized reward $\mathsf{R}_{v}$ is only known when vertex $v$ is visited. We assume that all reward distributions are discrete. Given a target $Q$, the goal is to find an (adaptive) tour originating from $r$ that collects a total reward at least $Q$ (or runs out of vertices), at the minimum expected length. A constant-factor approximation algorithm for stochastic $k$-TSP was obtained in [JLLS20], which was via a non-adaptive policy (so it is 1-threshold). Moreover, using a proof identical to that of Lemma 3.4, it follows that this problem is also up-monotone. Now, using Theorem 2.9, we obtain a polynomial time online learning algorithm for stochastic $k$-TSP having $O(1)$-regret $O(nD\sqrt{T\log T})$ where $D=\sum_{v\in V}d_{rv}$ corresponds to an upper bound on the (full) TSP cost.
Stochastic Submodular Maximization.
Consider a monotone submodular function $f:2^{E}\rightarrow\mathbb{Z}_{+}$ defined on a groundset $E$. Constrained submodular maximization involves selecting a subset $S\subseteq E$ of items in a downward-closed family ${\cal F}\subseteq 2^{E}$, that maximizes the function value $f(S)$. This is a fundamental problem in combinatorial optimization and good approximation algorithms are known for a variety of constraints ${\cal F}$ such as cardinality, matroids and knapsacks [NWF78, CCPV11, KST13, CVZ14]. We consider a stochastic variant, as in [AN16, GNS17, BSZ19], where each item $i\in E$ is active independently with some probability $p_{i}$ (and inactive otherwise). [^5] These papers also handle a more general setting where items have an arbitrary outcome space (not just binary). Moreover, the function value is accrued only on active items. The active/inactive status of any item is only known after selecting it. The goal now is to select a subset $S\in{\cal F}$ sequentially (and adaptively) so as to maximize the expected function value from the selected active items, i.e., $\mathbb{E}_{S,A}\left[f(S\cap A)\right]$ where $A\subseteq E$ is the random set of active items. Note that we can only select (and observe) a constrained subset of items, which corresponds to semi-bandit feedback. When ${\cal F}$ is a single matroid constraint, [AN16] obtained a $(1-\frac{1}{e}-\epsilon)$-approximation algorithm for the stochastic submodular maximization problem. Later, [GNS17, BSZ19] showed that a $\rho$-approximation algorithm for deterministic submodular maximization under any constraint ${\cal F}$ can be used to obtain a $\frac{\rho}{2}$-approximation algorithm for the stochastic problem under constraint ${\cal F}$. It is easy to see that this problem is down-monotone. So, Theorem 1.1 with $k=2$ (binary r.v.s) implies a $(1-\frac{1}{e}-\epsilon)$-regret of $O(nF\sqrt{T\log T})$ for stochastic submodular maximization over a matroid constraint, where $F$ is the maximum function value. Combined with the result of [BSZ19], we also obtain online learning algorithms for more general constraints ${\cal F}$.
Stochastic Submodular Cover
In the classic submodular cover problem [Wol82], there is a monotone submodular function $f:2^{E}\rightarrow\mathbb{Z}_{+}$ defined on a groundset $E$, where each item $i\in E$ has some cost $c_{i}$. We assume that the function is integer-valued with maximal value $Q$. The goal is to select a subset $S\subseteq E$ that “covers” the function, i.e., $f(S)=Q$. The objective is to minimize the total cost $\sum_{i\in S}c_{i}$ of the chosen items. Note that the submodular cover problem generalizes set cover. It is well-known that this problem can be approximated within a factor $1+\ln(Q)$ [Wol82], and no better approximation is possible [DS14]. We consider a stochastic version, where each item $i\in E$ is active independently with probability $p_{i}$, and the function value is accrued only on active items. The goal now is to select items sequentially (and adaptively) so as to cover function $f$ (or run out of items to select), at the minimum expected cost. This problem has been studied extensively [GK17, INZ16, HKP21], and a $1+\ln(Q)$ approximation algorithm is also known for it [HKP21]; this result even holds in the more general formulation with multiple outcomes for each item. Using the fact that the function $f$ is monotone, it is easy to see that stochastic submodular cover is up-monotone. So, using Theorem 1.1 with $k=2$, we obtain a polynomial time online learning algorithm for stochastic submodular cover having $(1+\ln Q)$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
Appendix
Appendix A Sampling-based Algorithms
A stronger monotonicity condition is that of strong monotonicity [GHTZ21], defined next.
Definition A.1 (Strong Monotonicity).
A stochastic problem is strongly up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{{E}}$, we have $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{D}}},\mathbf{{x}})]\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$, where $\sigma_{\mathbf{{D}}}$ is the optimal policy for instance $\mathcal{I}$ under distribution $\mathbf{{D}}$. The problem is said to be strongly down-monotone if $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{D}}},\mathbf{{x}})]\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ under the same conditions as above.
We note that strong monotonicity implies monotonicity. In prior work, [GHTZ21] gave optimal sample complexity bounds for stochastic optimization problems that exhibit the strong monotonicity condition. They also proved this property for problems including Prophet inequality and Pandora’s box. We make use of this property in §3, when we apply our result to these problems.
Below, we assume that $f_{max}=1$ by scaling. The sample-complexity result in [GHTZ21] is:
Theorem A.1 (Theorem 17 [GHTZ21]).
For any strongly monotone stochastic problem, suppose the number of samples is at least:
$C\cdot\frac{n}{\epsilon^{2}}\log\left(\frac{n}{\epsilon}\right)\log\left(\frac{nT}{\epsilon}\right)$
where $C>1$ is a sufficiently large constant. Then, there is an algorithm that gets an $\epsilon$-additive approximation to the optimum with probability at least $1-\frac{1}{T^{2}}$.
Below, we discuss what regret bounds can be achieved via this sampling-based approach, in order to compare to our results. In the full-feedback model, one can obtain regret bounds of $\sqrt{nT\log T}$, which is nearly optimal. However, for semi-bandit feedback that we consider (and for bandit feedback), such a “reduction” from sample-complexity bounds only provides a sub-optimal $\widetilde{O}(T^{2/3})$ regret. Therefore, we need new ideas to get the optimal $\widetilde{O}(\sqrt{T})$ regret bound in the semi-bandit model, which we do in this paper.
Full-feedback.
Recall that, in the full-feedback model, we get one sample from each r.v. in every time-step. We now describe the algorithm. The algorithm is straightforward: for each $t=1,2,\ldots T$, we use estimates from the prior $t-1$ time steps to obtain a policy to use for the $t^{\text{th}}$ time-step. See Algorithm 4 for a formal description.
Algorithm 4 Exploit
1:Input: one sample $\mathbf{{x}}=(x_{1},\cdots,x_{n})\sim\mathbf{{D}}$,
2:for $t=1,2,\ldots,T$ do
3:     let $\mathbf{{\widehat{D}_{t}}}\leftarrow$ frequentist estimate from previous $t-1$ time-steps
4:     run policy $\mathtt{OPT}(\mathbf{{\widehat{D}_{t}}})$ for the $t^{\text{th}}$ time-step
As a consequence of Theorem A.1, with probability at least $1-\frac{1}{T^{2}}$, we have that $\mathtt{OPT}(\mathbf{{\widehat{D}_{t}}})$ is an $\epsilon_{t}$-additive approximation to $\mathtt{OPT}(\mathbf{{D}})$ where $\epsilon_{t}\leq\sqrt{\frac{Cn\log(nT)}{t}}$. (To keep calculation simple, we ignore the $\log\frac{1}{\epsilon}$ dependence in Theorem A.1; so we are actually assuming a slightly stronger sample-complexity bound.) By union bound, this is true for all $t=1,\ldots,T$ with probability at least $1-\frac{1}{T}$. We consider this to be a good event, $G$. Under this event, the total regret, say $R_{T}$, can be bounded as follows.
$\mathbb{E}[R_{T}\mid G]\leq\sum_{t=1}^{T}\sqrt{\frac{{Cn\log\left(nT\right)}}{t}}=O\left(\sqrt{nT\log(nT)}\right).$
By law of total expectation we have
$\mathbb{E}[R_{T}]\leq\mathbb{E}[R_{T}\mid G]\cdot\mathbf{P}(G)+\mathbb{E}[R_{T}\mid\overline{G}]\cdot\mathbf{P}(\overline{G})\leq O\left(\sqrt{nT\log(nT)}\right)+T\cdot\frac{1}{T}=O\left(\sqrt{nT\log(nT)}\right).$
Semi-bandit feedback.
A significant challenge in the semi-bandit feedback model arises from our lack of control over the r.v.s from which we get samples. One potential strategy to address this issue is to artificially generate samples for an item $i$ by probing item $i$ first in the algorithm’s policy for that period. However, this approach comes with an inherent drawback – probing item $i$ first may result in a poor policy, and so we suffer a high regret in such periods. The standard explore-then-exploit algorithm that first gets $T^{2/3}$ samples for each r.v., and then plays the optimal policy (for the empirical distribution) for the remaining time steps only achieves regret $\widetilde{O}(T^{2/3})$.
In the hope of obtaining improved regret guarantees, we consider a different algorithm where we pair periods of exploration with some periods of exploitation, wherein we leverage the learned distribution. To do this, we divide the time horizon $T$ into meta periods, each of length $n+\beta n$ (we optimize $\beta$ later). In meta-period $h$, we first gather $n$ samples (one for each r.v.) and then use policy $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$ for $\beta n$ time-steps. Here $\mathbf{{\widehat{D}_{h}}}$ denotes the frequentist distribution obtained from the $h$ samples (until meta-period $h$) of $\mathbf{{D}}$. Note that we have no control over the r.v.s probed by $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$, and thus cannot use these samples in our analysis.
By Theorem A.1, with probability at least $1-\frac{1}{T^{2}}$, we have that $\mathtt{OPT}(\mathbf{{\widehat{D}_{h}}})$ is an $\epsilon_{h}$-additive approximation to $\mathtt{OPT}(\mathbf{{D}})$ where $\epsilon_{h}\leq\sqrt{\frac{Cn\log(nT)}{h}}$. By union bound, this is true for all meta-periods $h=0,1,\ldots,$ with probability at least $1-\frac{1}{T}$ (since $h\leq T$). As before, we consider this to be a good event, $G$. Under this event, the regret in meta-period $h$ is at most $n+\beta n\cdot\sqrt{\frac{Cn\log(nT)}{h}}$, and the total regret under $G$ is at most
$nH+C\beta n^{3/2}\sqrt{\log(nT)}\sum_{h=1}^{H}\sqrt{\frac{1}{h}}=nH+C\beta n^{3/2}\sqrt{H\log(nT)}$
where $H=\frac{T}{(1+\beta)n}$. The optimal choice for $\beta=O\left(\left(\frac{T}{n^{2}\log T}\right)^{1/3}\right)$, and so we again end up with a regret of $O\left((nT)^{2/3}(\log T)^{1/3}\right)$. To summarize, we are not aware of any generic approach that reduces sample-complexity bounds to regret minimization in the semi-bandit have $o(T^{2/3})$ regret.
Remark 1.
The sampling approach for semi-bandit feedback also applies to many problems (e.g., prophet inequality, Pandora’s box and series testing) in the bandit feedback model. Basically, for each item $i$ we need a policy where its objective corresponds to the value of r.v. $\mathsf{X}_{i}$. So, we can directly get $\widetilde{O}(T^{2/3})$ regret for these problems even with bandit feedback.
Remark 2.
We note that [GHTZ21] also give (slightly worse) sample complexity bounds for a broader class of problems that need not satisfy any monotonicity property, but have a finite support-size $k$. Using the approaches described above, we can convert these sample-complexity guarantees to obtain $O\left(\sqrt{nkT\log(nT)}\right)$ regret under full-feedback and $O\left((nkT)^{2/3}(\log T)^{1/3}\right)$ regret under semi-bandit (and bandit) feedback.